### v1

# Two algebraic proofs of the transcendence of $\mathrm{e}$ based on formal power series

Abstract
We first remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert. Then, using formal power series, we give two algebraic semiformal proofs of the same result. The first proof is a specialization of a proof of the Lindemann– Weierstrass theorem given by Beukers, Bézivin and Roba [2]. The second proof uses improper integrals of formal power series and is due to this author. At the beginning, we explain what “semiformal” means and indicate yet another analytical proof of the transcendence of $\mathrm{e}$. Like the two algebraic proofs, it avoids uncountable sets. It is due to this author in [4].
1 Hilbert’s proof that $\mathrm{e}$ is transcendental. Why this proof is not completely satisfactory. How to fix it
Hilbert’s proof [3] of the transcendence of the number $\mathrm{e}$ goes as follows. We suppose for the contradiction that there exists $n+1$, $n\in\mathbb{N}$ ($\{1,2,\dots\}$), integers $a_{0}$, $a_{1}$, $\dots$, $a_{n}$, not all of them 0, such that
$P:=a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0\,.$
We may assume that $a_{0}\neq 0$. For $r\in\mathbb{N}$ we define integral polynomials
$p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}\,$,
with $n$ as in $P$. We have the following expansions starting from the lowest powers of $x$:
$p_{r}(x)=(-1)^{n(r+1)}(n!)^{r+1}\cdot x^{r}+\bullet\cdot x^{r+1}+\dots$
and
$p_{r}(x+j)=\bullet\cdot x^{r+1}+\bullet\cdot x^{r+2}+\dots$
for $j=1,2,\dots,n$ and some $\bullet\in\mathbb{Z}$ (the integers). We consider integrals
$I_{r}=\int_{0}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x\,.$}
Using additivity of the improper integral $\int_{0}^{+\infty}$ we express the product $P\cdot I_{r}$ as
$0$ $=$ $P\cdot I_{r}=A_{r}+B_{r}$
$:=$ ${\textstyle\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x+\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x}\,.$
As for the quantity $A_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x$, it is easy to estimate it for $r\in\mathbb{N}$ by means of M-L bounds on integrals as
$|A_{r}|\leq c^{r}$
where $c>1$ is a constant depending only on the coefficients $a_{0}$, $a_{1}$, $\dots$, $a_{n}$.
The heart of the proof is the determination of the second quantity
$B_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x\,.$}
We claim that $B_{r}$ is an integral multiple of $r!$ and that $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. Since $c^{r}/r!\to 0$ as $r\to\infty$ for any constant $c>1$, the above equalities
$A_{r}+B_{r}=0,\ r\in\mathbb{N}\,$,
are contradictory and we are done. In order to prove the claim, we use the integral identity that
$\int_{0}^{+\infty}x^{n}\mathrm{e}^{-x}\,\mathrm{d}x=n!$}
for every $n\in\mathbb{N}_{0}$ ($\{0,1,\dots\}$). It is due to Euler and is easy to establish by induction on $n$ via integration by parts. By the linearity of the improper integral $\int_{0}^{+\infty}$, for any integers $b_{0}$, $b_{1}$, $\dots$, $b_{m}$ with $m\in\mathbb{N}_{0}$ we have more generally that
$\int_{0}^{+\infty}\big(\sum_{j=0}^{m}b_{j}x^{j}\big)\mathrm{e}^{-x}\,\mathrm{d}x=\sum_{j=0}^{m}b_{j}\cdot j!\,.$}
Then
$B_{r}$ $=$ ${\textstyle\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x}$
$=$ ${\textstyle\sum_{j=0}^{n}a_{j}\int_{j}^{+\infty}p_{r}((x-j)+j)\cdot\mathrm{e}^{-(x-j)}\,\mathrm{d}x}$
$=$ ${\textstyle\sum_{j=0}^{n}a_{j}\int_{0}^{+\infty}p_{r}(y+j)\cdot\mathrm{e}^{-y}\,\mathrm{d}y}$
$=$ $\pm a_{0}(n!)^{r+1}\cdot r!+m_{r}\cdot(r+1)!\,$,
for some $m_{r}\in\mathbb{Z}$. In the second line, we used the identity that $\mathrm{e}^{a+b}=\mathrm{e}^{a}\mathrm{e}^{b}$. In the third line, we changed the variable to y=xjy=x-j. In the fourth line, we used the above generalization of Euler’s identity and the expansions of $p_{r}(x+j)$, $j=0,1,\dots,n$, in lowest powers of $x$. It is clear that $B_{r}\in\mathbb{Z}$ and is divisible by $r!$. If $B_{r}=0$ then $(r+1)\,|\,a_{0}(n!)^{r+1}$. This divisibility is excluded when $r+1$ is coprime to the number $a_{0}\cdot n!$. Hence $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. $\Box$
Interestingly, Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof that uses only proper integrals $\int_{a}^{b}$.
Hilbert’s proof is elegant and short, but it is not completely satisfactory because it employs uncountable sets; such as, for example, the integrands in Euler’s identity. We explain how to fix it and how to prove the transcendence of the number $\mathrm{e}$ so that two conditions are satisfied.
Only hereditarily at most countable (HMC) sets are directly used in the proof.
No technical notion of mathematical logic such as, for example, the notion of a formula, is used in the proof.
A set $x$ is HMC if for every index $n\in\mathbb{N}_{0}$ and every chain of sets
$x_{n}\in x_{n-1}\in\dots\in x_{0}=x\,$,
the set $x_{n}$ is finite or countable. As for uses of mathematical logic, we quote a referee of [4]: “That basic real analysis is essentially a countable endeavor has been well-known for over 100 years. My secondhand understanding is that Weyl articulated this very well in Das Kontinuum, which nowadays we would say developed analysis in $\mathrm{ACA}_{0}$.” The reader can check [5] for $\mathrm{ACA}_{0}$. Our more demanding goal is to obtain a proof of the transcendence of $\mathrm{e}$ satisfying besides (A) also (B), that is, such that no HMC set and no mathematical logic is used.
There are (at least) three ways to achieve this goal. The first way is to replace real functions of the type $f\colon[a,b]\to\mathbb{R}$, where $a<b$ are real numbers and $\mathbb{R}$ is the set of real numbers, with the functions
$f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}\,$,
where $\mathbb{Q}$ is the (HMC) set of fractions. If real numbers are represented by Dedekind cuts, any such function $f$ is an HMC set. We realized the first way in [4]. Is there any difference in treatment of the two kinds of functions? The above mentioned referee thinks not: “Continuous functions on the reals are determined by their values on the rationals, so everything works out in the usual way. I don’t see how what the article is doing is so very different from developing analysis in (full) second-order arithmetic.” There is an obvious meta-argument refuting the referee’s first claim: if it were the case that “everything works out in the usual way,” then some enthusiasts would be teaching courses in analysis based on functions of the type $f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}$ for years. Since you have not heard about such courses, there must be some difficulties with this approach. Indeed, there are, but they can be overcome. The interested reader is referred to [4] for details. We include the referee’s second claim just to reiterate that our approach in [4] satisfies condition (B). In [4], we do not use technical notions of mathematical logic at all; therefore, our approach is very different from any countable treatment of real analysis offered by mathematical logic, such as, to give besides Das Kontinuum more modern references, [7] or Weaver’s forthcoming book [8].
We included this analytical digression to place the algebraic results of the present article in context. We present two algebraic proofs of the transcendence of $\mathrm{e}$ satisfying conditions (A) and (B). They are based on algebraic manipulations with formal power series. The first proof in Section 2 is a specialization of the proof in [2]. The second proof in Section 3 is due to this author. We presented it in early versions of [4], but we rework it here.
The word “directly” in condition (A) means that non-HMC sets may appear in the proof only by mention, as “harbors” of the main players. We may say, for example, that for every two real sequences $(a_{n}),(b_{n})\subset\mathbb{R}$ we have $\dots$; it is clear that $\mathbb{R}$ is used only indirectly and can be easily formulated away. We allow such uses of non-HMC sets for the sake of natural and easily understandable formulations.
In conclusion of this introduction, we review formal power series and explain the meaning of the adjective “semiformal”. If $R$ is an (integral) domain, we write $R[[x]]$ for the (usually non-HMC) set of maps $f\colon\mathbb{N}_{0}\to R$. They are traditionally written as infinite formal sums
$f=f(x)=\sum_{n\geq 0}a_{n}x^{n}$}\,,
where $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also by $x^{n}]f$. The set $R[[x]]$ is endowed with the operations
${\textstyle\sum_{n\geq 0}a_{n}x^{n}+\sum_{n\geq 0}a_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}(a_{n}+b_{n})x^{n}\,\text{ and}}$
${\textstyle\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}a_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}\,,}$
and neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. Then
R[[x]]=R[[x]], $\langle R[[x]],\,0,\,1$$,\cdot\rangle$
is a domain and $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is (formal) derivative,
$\big(\sum_{n\geq 0}a_{n}x^{n}\big)^{\prime}=\sum_{n\geq 0}(n+1)_{R}\cdot a_{n+1}x^{n}\,.$}
So for derivatives the characteristic of $R$ becomes relevant. Less often one uses the (formal) primitive
$\int\sum_{n\geq 0}a_{n}x^{n}=\sum_{n\geq 0}\frac{a_{n}}{n+1}x^{n+1}=\sum_{n\geq 1}\frac{a_{n-1}}{n}x^{n}\,$,}
but in this article, it is more important than the derivative. If $K$ is the normed field $K=\mathbb{R}$ or $K=\mathbb{C}$ (complex numbers), we denote by $K\{x\}$ the subdomain of $K[[x]]$ formed by convergent formal power series $f(x)=\sum_{n\geq 0}a_{n}x_{n}$ that satisfy
$\lim_{n\to\infty}a_{n}c^{n}=0$
for every $c\in K$. In other words, $f(x)$ has an infinite radius of convergence.
The binary and unary operations on $R[[x]]$ that we have listed so far, ++, $\cdot$, $(\dots)^{\prime}$, and $\int(\dots)$, are formal in the sense that, for example, $x_{n}](f\cdot g)$ is determined by finitely many coefficients in $f$ and $g$, and similarly for the other operations. In formal arguments involving elements $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$ it is also forbidden to use values
$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k}$}
of $f(x)$ on elements $c\in K$. It is still an open problem to find a fully formal proof of the transcendence of $\mathrm{e}$ based on formal power series; or at least no such proof is known to this author. However, in the next two sections we present two such proofs if formality is relaxed to quasiformality.
An argument involving elements of $K[[x]]$ is quasiformal if it uses only formal and quasiformal operations, and for every $f(x)\in K[[x]]$ appearing in it only at most countably many values $f(c)$ for $c\in K$ are used. Quasiformal operations allow uses of limit transitions in $K$ and of at most countably many values $f(c)$.
Quasiformal operations on $\mathbb{R}[[x]]$ are introduced, for example, in Definitions 3.3 and 3.11.
2 Specializing the proof of Beukers, Bézivin and Robba
The Lindemann–Weierstrass theorem says that
$b_{1}\mathrm{e}^{a_{1}}+b_{2}\mathrm{e}^{a_{2}}+\dots+b_{t}\mathrm{e}^{a_{t}}\neq 0$
whenever $b_{j}$ and $\alpha_{j}$ are $2t$, with $t\in\mathbb{N}$, algebraic numbers in $\mathbb{C}$ such that every $b_{j}\neq 0$ and the $\alpha_{j}$ are mutually distinct. See Baker [1, Theorem 1.4 on pp. 6–8] for an analytic proof. Using formal power series, Beukers, Bézivin and Robba gave in [2] another proof of the LW theorem. In this section, we specialize their proof to $b_{j},\alpha_{j}\in\mathbb{Z}$, which gives the transcendence of $\mathrm{e}$. In order to clearly describe logical structure of the proof, we completely reorganize the exposition in [2]. At the end of this section, we explain why this proof is semiformal (and not formal), and why it satisfies conditions (A) and (B).
We begin by recalling that a formal power series $f(x)$ in $\mathbb{C}[[x]]$ is rational if there exist two polynomials $p(x),q(x)\in\mathbb{C}[x]$ such that $q(0)=1$ and $q(x)f(x)=p(x)$; we may assume that $p(x)$ and $q(x)$ are coprime. If $f(x)$ and $g(x)$ in $\mathbb{C}[[x]]$ are rational, then so are $f(x)+g(x)$, $f(x)g(x)$, and $f^{\prime}(x)$. Any rational $f(x)\in\mathbb{C}[[x]]$ has a unique decomposition in partial fractions
$f(x)=r(x)+\sum_{j=1}^{t}\sum_{i=1}^{m_{j}}\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}\,$,}
where $r(x)\in\mathbb{C}[x]$, $t\in\mathbb{N}_{0}$, $m_{j}\in\mathbb{N}$, $\beta_{j,i},\alpha_{j}\in\mathbb{C}$ with $\beta_{j,m_{j}}\neq 0$ and the $\alpha_{j}$ are nonzero and mutually distinct; for $t=0$, the double sum is interpreted as 0, and the numbers $m_{j}$ and $\alpha_{j}$ come from the factorization
$q(x)=\prod_{j=1}^{t}(1-\alpha_{j}x)^{m_{j}}\,.$}
The partial fractions are the formal power series
$\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}=\beta_{j,i}\sum_{n\geq 0}\binom{-i}{n}\alpha_{j}^{n}x^{n}\,.$}
We say that $1/\alpha_{j}$ is a pole of $f(x)$ with order $m_{j}$.
Proposition 2.1
If $p(x)\in\mathbb{C}[x]$, $m\in\mathbb{N}_{0}$, $c\in\mathbb{C}^{*}$ ($\mathbb{C}\setminus\{0\}$) and if $A(x)$ in $\mathbb{C}[[x]]$ is rational, then the rational formal power series
$p(x)A(x)+cx^{m}A^{\prime}(x)$
either has no pole (is a polynomial) or all its poles have orders at least $2$.
Proof. This follows from the definition of the decomposition of a rational formal power series in partial fractions. The product $p(x)A(x)$ possibly only decreases the orders of the poles of $A(x)$, or cancels them and does not create any new poles. The product $cx^{m}A^{\prime}(x)$ preserves all poles of $A(x)$, increases their orders by $1$, and does not create any new poles. The sum $p(x)A(x)+cx^{m}A^{\prime}(x)$ merges the two sets of poles. Therefore, the resulting set of poles with their orders is that of $cx^{m}A^{\prime}(x)$. $\Box$
We prove that $\mathrm{e}$ is transcendental by means of the following theorem on rationality of formal power series, which is extracted from the arguments in [2].
Theorem 2.2
Let $t\in\mathbb{N}$, $b_{1}$, $b_{2}$, $\dots$, $b_{t}$ be nonzero integers and $\alpha_{1}$, $\alpha_{2}$, $\dots$, $\alpha_{t}$ be distinct natural numbers. For every $n\in\mathbb{N}_{0}$ we define integers
$u_{n}=b_{1}\alpha_{1}^{n}+\dots+b_{t}\alpha_{t}^{n}\,\text{ and }\,v_{n}=n!\cdot{\textstyle\sum_{r=0}^{n}u_{r}/r!}\,.$
Suppose that for some constant $A\geq 1$ we have bound
$v_{n}=O(A^{n})\,\text{ for }\,n\in\mathbb{N}_{0}\,.$
Then the formal power series $\sum_{n\geq 0}v_{n}x^{n}$ is rational, because
$\big(1-a_{1}x-a_{2}x^{2}-\dots-a_{t}x^{t}\big)^{k}\sum_{n\geq 0}v_{n}x^{n}\in\mathbb{Z}[x]$}
for some $k\in\mathbb{N}$, where the integers $a_{j}$ are the coefficients in
$\prod_{j=1}^{t}(1-\alpha_{j}x)=1-a_{1}x-a_{2}x^{2}-\dots-a_{t}x^{t}\,.$}
Proof. For $k,n\in\mathbb{N}_{0}$ we define integers $v_{n}(k)$ by the relations
$\sum_{n\geq 0}v_{n}(k)x^{n}=\big(1-a_{1}x-a_{2}x^{2}-\dots-a_{t}x^{t}\big)^{k}\sum_{n\geq 0}v_{n}x^{n}\,.$}
Then for every $n\geq t$ and $k\in\mathbb{N}_{0}$ we have the recurrence
$v_{n}(k+1)=v_{n}(k)-a_{1}v_{n-1}(k)-\dots-a_{t}v_{n-t}(k)\,.$
We establish for $n,k\in\mathbb{N}_{0}$ with $n\geq tk$ two properties of the numbers $v_{n}(k)$.
  1. 1.
    We have bound $|v_{n}(k)|\leq cA^{n}C^{k}$ where $C=1+|a_{1}|+\dots+|a_{t}|$ and $c>0$ is a constant.
  2. 2.
    $k!\,|\,v_{n}(k)$.
The bound in item 1 follows by induction on $k$ and the assumed bound on $v_{n}=v_{n}(0)$. For $k=0$ the bound clearly holds. Let $n,k\in\mathbb{N}_{0}$ with $n\geq t(k+1)$. Then $n-t\geq tk$, and by the above recurrence and induction we have
$|v_{n}(k+1)|$ $=$ $|v_{n}(k)-a_{1}v_{n-1}(k)-\dots-a_{t}v_{n-t}(k)|$
$\leq$ $c\big(A^{n}C^{k}+A^{n-1}C^{k}|a_{1}|+\dots+A^{n-t}C^{k}|a_{t}|\big)\leq cA^{n}C^{k+1}\,.$
In order to prove the divisibility in item 2, we note that $v_{0}=u_{0}$ and that for $n\in\mathbb{N}$,
$\frac{v_{n}}{n!}-\frac{v_{n-1}}{(n-1)!}=\frac{u_{n}}{n!}\,\text{ (hence $v_{n}-nv_{n-1}=u_{n}$)}\,.$}
For $k=0$ the divisibility trivially holds. So let $k\in\mathbb{N}$. We set $(0)_{0}=(n)_{0}=1$ and $(n)_{m}=n(n-1)\dots(n-m+1)$ for $m\in\mathbb{N}$. For $n\in\mathbb{N}_{0}$ we combine $\min(n,k)$ displayed differences, multiply the result by $n!$, and get that
$v_{n}=(n)_{0}u_{n}+(n)_{1}u_{n-1}+\dots+(n)_{k-1}u_{n-k+1}+(n)_{k}v_{n-k}$}\,,
where the $u_{\dots}$ and $v_{\dots}$ with negative indices are defined arbitrarily, for example as 0. We denote the last summand $(n)_{k}v_{n-k}$ ($\in\mathbb{Z}$) by $w_{n}$. Note that always $k!\,|\,w_{n}$ because $w_{n}=k!\cdot\binom{n}{k}v_{n-k}$ for $n\geq k$, and $w_{n}=0$ for $0\leq n<k$. The difference of the formal power series $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ is
$V(x)-W(x)=\sum_{n\geq 0}\big((n)_{0}u_{n}+(n)_{1}u_{n-1}+\dots+(n)_{k-1}u_{n-k+1}\big)x^{n}\,.$}
We claim that
$V(x)-W(x)=\frac{p(x)}{(1-a_{1}x-\dots-a_{t}x^{t})^{k}}=:\frac{p(x)}{q(x)}$}
for some polynomial $p(x)\in\mathbb{Z}[x]$ with degree less than $tk$. Then for every $n\geq tk$ we have
$v_{n}(k)=[x^{n}]q(x)V(x)=[x^{n}]q(x)W(x)$
and $k!\,|\,v_{n}(k)$ because every coefficient in $W(x)$ is divisible by $k!$. We prove the claim: for any $r$ with $0\leq r\leq k-1$ we indeed have, as $\binom{n}{r}=(-1)^{n-r}\binom{-r-1}{n-r}$ for $n\geq r$, that
${\textstyle\sum_{n\geq 0}(n)_{r}u_{n-r}x^{n}}$ $=$ ${\textstyle\sum_{j=1}^{t}r!\cdot b_{j}x^{r}\sum_{n\geq r}\binom{n}{r}\alpha_{j}^{n-r}x^{n-r}}$
$=$ ${\textstyle\sum_{j=1}^{t}r!\cdot b_{j}x^{r}(1-\alpha_{j}x)^{-r-1}=\frac{p_{r}(x)}{(1-a_{1}x-\dots-a_{t}x^{t})^{r+1}}}$
for some polynomial $p_{r}(x)\in\mathbb{Z}[x]$ with degree less than $t(r+1)$.
Now the two properties of numbers $v_{n}(k)$ imply that if $n\geq tk$ and $v_{n}(k)\neq 0$, then
$k!\leq|v_{n}(k)|\leq cA^{n}C^{k}\,.$
In other words, if $n\geq tk$ and if $k!>cA^{n}C^{k}$, then $v_{n}(k)=0$. Using this implication and the modification
$v_{n}(k)=v_{n}(k+1)+a_{1}v_{n-1}(k)+\dots+a_{t}v_{n-t}(k),\ n\geq t\,$,
of the above recurrence, we show that for a fixed $k\in\mathbb{N}$ we have $v_{n}(k)=0$ for every large $n$. This will prove that the formal power series $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ is rational. So let $k_{0}\in\mathbb{N}$ be such that $k!>cA^{2tk}C^{k}$ for every $k\geq k_{0}$, and let
$N$ $=$ $\{\langle k,\,n\rangle\in\mathbb{N}^{2}\colon\;k\geq k_{0}\wedge tk\leq n\leq 2tk\}\,\text{ and}$
$M$ $=$ $\{\langle k,\,n\rangle\in\mathbb{N}^{2}\colon\;k\geq k_{0}\wedge n\geq 2tk\}\,.$
Thus $\langle k,n\rangle\in N$ $\Rightarrow$ $v_{n}(k)=0$. For $\langle k,n\rangle\in\mathbb{N}^{2}$ we consider the norm $\|\langle k,n\rangle\|=n-2tk$ ($\in\mathbb{Z}$). We name the ordered pairs in the modified recurrence:
$\gamma=\langle k,\,n\rangle,\,\beta_{0}=\langle k+1,\,n\rangle,\,\beta_{1}=\langle k,\,n-1\rangle,\,\dots,\,\beta_{t}=\langle k,\,n-t\rangle\,.$
The following three properties of the norm $\|\dots\|$ and the pairs $\gamma$, $\beta_{0}$, $\dots$, $\beta_{t}$ are easy to prove.
  1. 1.
    The norm $\|\dots\|$ attains a minimum value on the set $M$ (namely, zero).
  2. 2.
    If $\gamma\in M$ then $\beta_{i}\in M\cup N$ for every $i=0,1,\dots,t$.
  3. 3.
    For every $i=0,1,\dots,t$, the norm $\|\beta_{i}\|<\|\gamma\|$.
From these three properties and from the modified recurrence, it follows by induction on the norm that $v_{n}(k)=0$ for every $\langle k,n\rangle\in M$. This concludes the proof because we see that every $k\geq k_{0}$ has the required property. $\Box$
We prove, with the help of Proposition 2.1 and Theorem 2.2, that the number $\mathrm{e}$ is transcendental. We assume, for the contradiction, that $b_{j},\alpha_{j}\in\mathbb{Z}$, $j\in[t]$ ($\{1,2,\dots,t\}$), are $2t$ integers such that every $b_{j}\neq 0$, the $\alpha_{j}$ are mutually distinct, and that
$\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}}=0\,.$}
We may assume that $\alpha_{j}\in\mathbb{N}$. We set $A:=\max(\{\alpha_{1},\dots,\alpha_{t}\})$ and define $u_{n}$ and $v_{n}$ as in Theorem 2.2. Then
$\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\sum_{n\geq 0}\frac{\alpha_{j}^{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}x}$}
and
$\sum_{n\geq 0}\frac{u_{n}}{n!}=\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}}=0\,.$}
From the definition of $u_{n}$ and $A$ it easily follows that $u_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$. Since by the definition of $v_{n}$ we have
$|v_{n}|={\textstyle n!\cdot\big|\sum_{r=0}^{n}u_{r}/r!\big|=n!\cdot\big|\sum_{r=n+1}^{\infty}u_{r}/r!\big|}$
— now we invoke the contradictory assumption — we get the bound
$|v_{n}|$ $\leq$ ${\textstyle\frac{|u_{n+1}|}{n+1}+\frac{|u_{n+2}|}{(n+1)(n+2)}+\dots\leq cA^{n}\big(\frac{A}{n+1}+\frac{A^{2}}{(n+1)(n+2)}+\dots\big)}$
$\leq$ ${\textstyle cA^{n}\sum_{j\geq 0}\frac{A^{j}}{j!}=c\,\mathrm{e}^{A}\cdot A^{n}\,,}$
where $c>0$ is an absolute constant. Thus $v_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$. Let $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ ($\in\mathbb{Z}[[x]]$). As we know from the previous proof, $v_{n}-nv_{n-1}=u_{n}$ for $n\in\mathbb{N}$ and $v_{0}=u_{0}$. In terms of formal power series we get
$(1-x)V(x)-x^{2}V^{\prime}(x)=\sum_{n\geq 0}(v_{n}-nv_{n-1})x^{n}=\sum_{n\geq 0}u_{n}x^{n}=\sum_{j=1}^{t}\frac{b_{j}}{1-\alpha_{j}x}$}
(with $v_{-1}:=0$). However, the equality
$(1-x)V(x)-x^{2}V^{\prime}(x)=\sum_{j=1}^{t}\frac{b_{j}}{1-\alpha_{j}x}$}
contradicts Proposition 2.1 because by Theorem 2.2 the formal power series $V(x)$ is rational. $\Box$
We reiterate that the proof in [2] is more general as it establishes the full LW theorem. In this situation, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of. The previous proof works with formal power series, but it is only semiformal and not formal because the key player is the infinite sum
$\sum_{n\geq 0}\frac{u_{n}}{n!}=0$}
which is the value $f(1)$ of the formal power series $f(x)=\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}$. In fact, this is the only non-formal feature of it.
It is clear that the previous proof satisfies condition (A), which states that all directly used sets are HMC. It works directly only with complex formal power series, which are maps from $\mathbb{N}_{0}$ to $\mathbb{C}$. As explained in Section 1, non-HMC sets like $\mathbb{C}[[x]]$ are not directly used in the proof. Condition (B), which states that no formalism of mathematical logic is used, is clearly satisfied as well. Although the article [2] of Beukers, Bézivin, and Robba is a pioneering one, as it is the first proof of the transcendence of the number $\mathrm{e}$ (in fact, of course, of the full LW theorem) that satisfies conditions (A) and (B), it is little known; in the database of Mathematical Reviews, it currently has only 11 citation (January 2026).
3 The proof by formal integration
We present the second semiformal proof of the transcendence of $\mathrm{e}$. It is due to this author and first appeared in early versions of [4]. The idea is simply to translate Hilbert’s proof, which we described at the beginning, into the language of formal power series. At the end of the section we explain why the proof is semiformal and why conditions (A) and (B) are satisfied.
We work in the domains of formal power series $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$. We introduce additional operations on $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$ besides those we reviewed in Section 1. For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
$f(ax)=\sum_{n\geq 0}a^{n}a_{n}x^{n}\,.$}
This is a formal operation. Clearly, for every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
$(af+bg)(cx)=af(cx)+bg(cx)\,\text{ and }\,(fg)(cx)=f(cx)g(cx)\,.$
If $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ then $f(0x)=a_{0}$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$.
In Section 1 we defined the derivative $f^{\prime}$ and the primitive $\int f$ of a formal power series $f\in\mathbb{R}[[x]]$. In the next proposition, whose straightforward proof we omit, we collect properties of these operations.
Proposition 3.1
Let $f,g\in\mathbb{R}[[x]]$ and $a,b\in\mathbb{R}$. The following holds.
  1. 1.
    $(af+bg)^{\prime}=af^{\prime}+bg^{\prime}$, $(fg)^{\prime}=f^{\prime}g+fg^{\prime}$ and $f(ax)^{\prime}=af^{\prime}(ax)$.
  2. 2.
    $\int(af+bg)=a\int f+b\int g$ and $\int f(ax)=\frac{1}{a}(\int f)(ax)$ if $a\neq 0$.
  3. 3.
    If $f\in\mathbb{R}\{x\}$ then $f^{\prime},\int f\in\mathbb{R}\{x\}$.
  4. 4.
    $\big(\int f\big)^{\prime}=f$ and $\int(f^{\prime})=f-[x^{0}]f$.
We define the (formal) exponential.
Definition 3.2
The formal power series
$\mathrm{Exp}(x)=\sum_{n\geq 0}\frac{1}{n!}x^{n}=1+x+\frac{1}{2}x^{2}+\frac{1}{6}x^{3}+\dots\ \ (\in\mathbb{Q}[[x]])$}
is called the (formal) exponential.
It is easy to show that $\mathrm{Exp}\in\mathbb{R}\{x\}$, $\mathrm{Exp}^{\prime}=\mathrm{Exp}$ and that $\int\mathrm{Exp}=\mathrm{Exp}-1$.
We introduce a semiformal operation.
Definition 3.3 (semiformal shifts)
Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ be in $\mathbb{R}\{x\}$ and let $b\in\mathbb{R}$. We define the formal power series $f(x+b)$ in $\mathbb{R}[[x]]$ by
$f(x+b)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)x^{n}\;.$}
The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent, and $f(x+b)\in\mathbb{R}\{x\}$. It is easy to see that $f(x+0)=f(x)$ and that if $f(x)\in\mathbb{R}[x]$ then $f(x+b)$ is the usual substitution in a polynomial. In the next two propositions we show that semiformal shifts behave in the expected way.
Proposition 3.4
Let $f(x)\in\mathbb{R}\{x\}$, $c,d\in\mathbb{R}$ and $g(x)=f(x+c)$. Then
g(x+d)=f(x+(c+d)).g(x+d)=f(x+(c+d))\;.
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $c$, $d$ and $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ be as stated. Then
$b_{n}=\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,g(x+d)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}\big)x^{n}\,.$}
Thus the coefficient of $x^{n}$ in $g(x+d)$ equals
${\textstyle\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}=\sum_{m\geq n}\binom{m}{n}\big(\sum_{l\geq m}\binom{l}{m}a_{l}c^{l-m}\big)d^{m-n}=}$
$={\textstyle\sum_{m\geq n}\sum_{l\geq m}\frac{l!}{n!(m-n)!(l-m)!}a_{l}c^{l-m}d^{m-n}=}$
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{m=n}^{l}\binom{l-n}{l-m}c^{l-m}d^{m-n}=}$
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{j=0}^{l-n}\binom{l-n}{j}c^{j}d^{l-n-j}=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}\,,}$
which is the coefficient of $x^{n}$ in $f(x+(c+d))$. In the third line, we changed the order of summation. This is allowed because $f(x)$ is convergent. In the fourth line, we used the binomial theorem. $\Box$
Proposition 3.5
Let $f,g\in\mathbb{R}\{x\}$ and $c\in\mathbb{R}$. Then
$(f\cdot g)(x+c)=f(x+c)\cdot g(x+c)\;.$
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ and $c$ be as stated. The coefficient of $x^{n}$ in the formal power series $f(x+c)$, respectively $g(x+c)$, equals $\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}$, respectively $\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}$. Thus the coefficient of $x^{n}$ in the product $f(x+c)\cdot g(x+c)$ equals
${\textstyle\sum_{k=0}^{n}\big(\sum_{m\geq k}\binom{m}{k}a_{m}c^{m-k}\cdot\sum_{l\geq n-k}\binom{l}{n-k}b_{l}c^{l+k-n}\big)=}$
${\textstyle=\sum_{k=0}^{n}\sum_{m\geq k}\sum_{l\geq n-k}\binom{m}{k}\binom{l}{n-k}a_{m}b_{l}c^{l+m-n}=}$
${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\sum_{k=0}^{n}\binom{m}{k}\binom{l}{n-k}a_{m}b_{l}c^{l+m-n}=}$
${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\binom{m+l}{n}a_{m}b_{l}c^{l+m-n}=\sum_{j\geq n}\binom{j}{n}\big(\sum_{m=0}^{j}a_{m}b_{j-m}\big)c^{j-n}\,,}$
which is the coefficient of $x^{n}$ in $(f\cdot g)(x+c)$. In the third line, we changed the order of summation. This is allowed since $f(x)$ is convergent. In the fourth line, we used Vandermonde’s identity ([6]). $\Box$
We prove a semiformal exponential identity.
Theorem 3.6
Let $a,b\in\mathbb{R}$ with $a\neq 0$ and $f(x)=\mathrm{Exp}\big(x+\frac{b}{a}\big)$. Then
$f(ax)=\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$
Proof. Let $a$, $b$ and $f(x)$ be as stated. Then, indeed,
$f(ax)$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}(a^{m}/m!)(b/a)^{m-n}\big)x^{n}}$
$=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}=\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)\frac{a^{n}}{n!}x^{n}}$
$=$ $\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\;.$
$\Box$
We introduce the Newton integral of convergent formal power series.
Definition 3.7 (Newton integral of fps)
Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}\in\mathbb{R}\{x\}$ and let $u,v\in\mathbb{R}$. We define the Newton integral of $f(x)$ from $u$ to $v$ as the difference
${\textstyle\int_{u}^{v}f=\int_{u}^{v}f(x)}$ $:=$ ${\textstyle\big(\int f\big)(v)-\big(\int f\big)(u)}$
$=$ ${\textstyle\sum_{n\geq 0}a_{n}\cdot\frac{v^{n+1}}{n+1}-\sum_{n\geq 0}a_{n}\cdot\frac{u^{n+1}}{n+1}}$
of the values of the formal primitive of $f$ at the numbers $v$ and $u$.
The integral is always defined by item 3 of Proposition 3.1 because $\int f\in\mathbb{R}\{x\}$. The Newton integral is a semiformal operation because it uses two values of a formal power series. We establish for it three integral transformations: linear combination, additive combination, and shift of the integration interval. The first two are quite straightforward.
Proposition 3.8
Let $f,g\in\mathbb{R}\{x\}$ and $u,v,a,b\in\mathbb{R}$. Then
$\int_{u}^{v}(af+bg)=a\int_{u}^{v}f+b\int_{u}^{v}g\,.$}
Proof. This follows from the equality $\int(af+bg)=a\int f+b\int g$ in Proposition 3.1. $\Box$
Proposition 3.9
Let $f\in\mathbb{R}\{x\}$ and $u,v,w\in\mathbb{R}$. Then
$\int_{u}^{w}f=\int_{u}^{v}f+\int_{v}^{w}f\,.$}
Proof. Indeed,
$\big(\int f\big)(w)-\big(\int f\big)(u)=\big(\int f\big)(v)-\big(\int f\big)(u)+\big(\int f\big)(w)-\big(\int f\big)(v)\,.$}
$\Box$
The third transformation needs some work.
Proposition 3.10
Let $f\in\mathbb{R}\{x\}$ and $u,v,b\in\mathbb{R}$. Then
$\int_{u}^{v}f(x+b)=\int_{u+b}^{v+b}f(x)\,.$}
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $u$, $v$ and $b$ be as stated. Then
${\textstyle\int_{u}^{v}f(x+b)=}$
${\textstyle=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)\frac{v^{n+1}}{n+1}-\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)\frac{u^{n+1}}{n+1}}$
$={\textstyle\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{n=0}^{m}\binom{m+1}{n+1}b^{m-n}v^{n+1}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{n=0}^{m}\binom{m+1}{n+1}b^{m-n}u^{n+1}}$
${\textstyle=\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{l=0}^{m+1}\binom{m+1}{l}b^{m+1-l}v^{l}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{l=0}^{m+1}\binom{m+1}{l}b^{m+1-l}u^{l}}$
${\textstyle=\sum_{m\geq 0}a_{m}\frac{(v+b)^{m+1}}{m+1}-\sum_{m\geq 0}a_{m}\frac{(u+b)^{m+1}}{m+1}=\int_{u+b}^{v+b}f(x)\,.}$
In the third line, we used the identity $\binom{m+1}{n+1}=\frac{m+1}{n+1}\binom{m}{n}$ and we exchanged the order of summation. This is allowed since the involved series absolutely converge. In the fourth line, we add and subtract the term $b^{m+1}$ in the two inner sums. In the fifth line, we use the binomial theorem. $\Box$
We introduce the improper Newton integral of formal power series.
Definition 3.11 (improper Newton $\int$ of fps)
Let $u\in\mathbb{R}$ and $f(x)\in\mathbb{R}\{x\}$. If for every real sequence $(b_{n})$ going to $\infty$ the limit
$L=\lim_{n\to\infty}\int_{u}^{b_{n}}f\ \ (\in\mathbb{R})$}
exists, then $L$ does not depend on the sequence $(b_{n})$ and we define
$\int_{u}^{+\infty}f=\int_{u}^{+\infty}f(x):=L\,.$}
We call this integral the improper Newton integral from $u$ to $\infty$ of the formal power series $f(x)$.
We leave the proof of the independence of the limit $L$ on the sequences $(b_{n})$ to the reader as an exercise.
We obtain for the improper integral the same three integral transformations as before.
Proposition 3.12
Let $f,g\in\mathbb{R}\{x\}$ and let $u,a,b\in\mathbb{R}$. Then the equality
$\int_{u}^{+\infty}(af+bg)=a\int_{u}^{+\infty}f+b\int_{u}^{+\infty}g$}
holds whenever the last two improper Newton integrals exist.
Proof. This follows from Proposition 3.8 and the arithmetic of limits of real sequences. $\Box$
Proposition 3.13
Let $f\in\mathbb{R}\{x\}$ and $u,v\in\mathbb{R}$. The equality
$\int_{u}^{+\infty}f=\int_{u}^{v}f+\int_{v}^{+\infty}f$}
holds whenever one of the two improper Newton integrals exists.
Proof. Let $f$, $u$ and $v$ be as stated, let $(b_{n})$ be any real sequence going to $\infty$ and let us suppose that the former improper integral exists. Thus the limit $L=\lim_{n\to\infty}\big(\int f\big)(b_{n})$ exists. But then
${\textstyle\int_{u}^{+\infty}f=}$
${\textstyle=L-\big(\int f\big)(u)=\big(\int f\big)(v)-\big(\int f\big)(u)+L-\big(\int f\big)(v)=}$
${\textstyle=\int_{u}^{v}f+\int_{v}^{+\infty}f\,.}$
With the assumption of existence of $\int_{v}^{+\infty}f$ we argue similarly. $\Box$
The third transformation is again less immediate.
Proposition 3.14
Let $f\in\mathbb{R}\{x\}$ and $u,b\in\mathbb{R}$. The equality
$\int_{u}^{+\infty}f(x+b)=\int_{u+b}^{+\infty}f(x)$}
holds whenever one of the two improper Newton integrals exists.
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $u$ and $b$ be as stated. Let $(c_{n})$ be any real sequence converging to $\infty$, and let $L$ and $R$ be the integrals on the left-hand side and the right-hand side, respectively. It suffices to show the implication $\exists\,R\Rightarrow R=L$; the opposite implication then follows. Indeed, we suppose that integral $L$ exists and write it as
$L=\int_{u+b-b}^{+\infty}f(x+b)\,.$}
We set $g(x)=f(x+b)$, use the assumed implication with b-b and get that
[…]
Thus, the implication $\exists\,L\Rightarrow L=R$ holds as well.
We suppose that integral $R$ exists. We replace the test sequence $(c_{n})$ with $(c_{n}+b)$ and get
$R$ $=$ $\lim_{m\to\infty}{\textstyle\big(\sum_{n\geq 0}a_{n}\frac{(c_{m}+b)^{n+1}}{n+1}\big)-\sum_{n\geq 0}a_{n}\frac{(u+b)^{n+1}}{n+1}}$
$=$ $\lim_{m\to\infty}{\textstyle\big(\sum_{l\geq 0}\sum_{n\geq l}\binom{n+1}{l+1}a_{n}b^{n-l}\frac{c_{m}^{l+1}}{n+1}\big)-\sum_{l\geq 0}\sum_{n\geq l}\binom{n+1}{l+1}a_{n}b^{n-l}\frac{u^{l+1}}{n+1}}$
$=$ $\lim_{m\to\infty}{\textstyle\big(\sum_{l\geq 0}\sum_{n\geq l}\binom{n}{l}a_{n}b^{n-l}\frac{c_{m}^{l+1}}{l+1}\big)-\sum_{l\geq 0}\sum_{n\geq l}\binom{n}{l}a_{n}b^{n-l}\frac{u^{l+1}}{l+1}}$
$=$ $\lim_{m\to\infty}{\textstyle\big(\int f(x+b)\big)(c_{m})}-{\textstyle\big(\int f(x+b)\big)(u)}=L\,.$
In the second line, we used the binomial theorem, and we changed the order of summation, which is allowed by the convergence of $f(x)$. Also, we canceled the summands corresponding to $l+1=0$. In the third line, we used the transformation $\binom{l+1}{n+1}=\frac{l+1}{n+1}\binom{l}{n}$. In the fourth line, we use the definition of formal primitive and Definition 3.3. $\Box$
We need bounds on integrals
$\int_{0}^{i}p(x)\mathrm{Exp}(-x)$}
where $i\in\mathbb{N}_{0}$, $p(x)$ is an integral polynomial and $\mathrm{Exp}(-x)=\mathrm{Exp}((-1)x)$. In the semiformal approach, this appears problematic because one deduces the standard LM bound by considering all values of the integrand on $[0,i]$ or $0,i]\cap\mathbb{Q}$. Instead, we use the next substitute LM bound, which can be proven by means of just two values of the primitive.
Proposition 3.15
For every $i,k\in\mathbb{N}_{0}$,
$\big|\int_{0}^{i}x^{k}\cdot\mathrm{Exp}(-x)\big|\leq i^{k+1}\mathrm{e}^{i}\,.$}
Proof. The integrand
$\text{$\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!}x^{n+k}$ has the primitive $\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!(n+k+1)}x^{n+k+1}$}\,.$}
Thus the Newton integral is in absolute value at most $\sum_{n=0}^{\infty}\frac{1}{n!}i^{n+k+1}+0=i^{k+1}\mathrm{e}^{i}$. $\Box$
It is important that the bound is exponential and not factorial in $k$.
The key step in Hilbert’s proof of the transcendence of $\mathrm{e}$ was Euler’s identity. Now we prove it in the setting of the improper Newton integral of formal power series.
Theorem 3.16
For every $k\in\mathbb{N}_{0}$, the integral
$I_{k}=\int_{0}^{+\infty}x^{k}\cdot\mathrm{Exp}(-x)=\int_{0}^{+\infty}\sum_{n=k}^{\infty}\frac{(-1)^{n-k}}{(n-k)!}x^{n}=k!\,.$}
Proof. We proceed by induction on $k$. Let $(a_{n})$ be any real sequence going to $\infty$. For $k=0$ we use that
$\int\mathrm{Exp}(-x)=\sum_{n\geq 0}\frac{(-1)^{n}}{(n+1)!}x^{n+1}=1-\mathrm{Exp}(-x)\,.$}
Hence
$I_{0}=\lim_{n\to\infty}{\textstyle(1-\mathrm{Exp}(-a_{n}))-(1-\mathrm{Exp}(-0))=1-0=1\,.}$
For $k\in\mathbb{N}_{0}$, let $F_{k}=x^{k}\cdot\mathrm{Exp}(-x)$. Let $k\in\mathbb{N}$. By item 1 of Proposition 3.1 and the derivative of $\mathrm{Exp}(-x)$,
$F_{k}^{\prime}=\big(x^{k}\cdot\mathrm{Exp}(-x)\big)^{\prime}=kx^{k-1}\cdot\mathrm{Exp}(-x)-x^{k}\cdot\mathrm{Exp}(-x)=k\cdot F_{k-1}-F_{k}\,.$
By item 4 of Proposition 3.1, $\int F_{k}^{\prime}=F_{k}$. Thus by Proposition 3.12 and induction,
$I_{k}$ $=$ $\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k}=k\cdot}\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k-1}-}\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k}^{\prime}}$
$=$ $k\cdot I_{k-1}-\lim_{n\to\infty}F_{k}(a_{n})+F_{k}(0)$
$=$ $k\cdot(k-1)!-0+0=k!\;.$
$\Box$
We are ready to translate Hilbert’s proof of the transcendence of $\mathrm{e}$ into formal power series. We again assume for the contradiction that
$a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0$
where $a_{i}\in\mathbb{Z}$, $n\in\mathbb{N}$, and $a_{0}\neq 0$. We multiply this equality by the improper Newton integral (of formal power series)
$\int_{0}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)\,$,}
where $p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$ (as in Section 1). This integral exists by Proposition 3.12 and Theorem 3.16. Using Proposition 3.13 we get that
$0$ $=$ $A_{r}+B_{r}$
$:=$ ${\textstyle\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\cdot\int_{0}^{i}p_{r}(x)\cdot\mathrm{Exp}(-x)+\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\cdot\int_{i}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)}$
for every $r\in\mathbb{N}$. We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded.
The determination of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that we calculate semiformally. We define abbreviations
$\mathrm{Exp}(-(x-i))$ $=$ $f((-1)x),\ f(x)=\mathrm{Exp}(x+(-i))\,\text{ and}$
$\mathrm{Exp}(-((x-i)+i))$ $=$ $f((-1)x),\ f(x)=g(x+i),\ g(x)=\mathrm{Exp}(x+(-i))\,.$
Then
$B_{r}$ $=$ ${\textstyle\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{i}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)}$
[…]
Recall that $x$ is a formal variable. Using Proposition 3.12, Theorem 3.16, and the definition of the polynomial $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
It remains to bound the numbers
$A_{r}=\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{0}^{i}p_{r}(x)\cdot\mathrm{Exp}(-x)\,.$}
The polynomial
$p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$
has degree less than $(n+1)(r+1)$ and each coefficient of it is in absolute value at most $l^{r+1}$ where $l\in\mathbb{N}$ depends only on $n$. By Propositions 3.8 and 3.15 each of the above $n+1$ Newton integrals in $A_{r}$ is in absolute value at most
$(n+1)(r+1)\cdot l^{r+1}\cdot n^{(n+1)(r+1)}\cdot\mathrm{e}^{n}\,.$
Thus $|A_{r}|\leq c^{r}$ for every $r\in\mathbb{N}$ and some constant $c\geq 1$. We get the same contradiction as in Section 1. $\Box$
The previous proof is quasiformal and not formal because of Definitions 3.3 and 3.11. It clearly satisfies conditions (A) and (B).

### v3

@@
 Abstract
-We first remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert.
-Then, using formal power series, we give two algebraic semiformal proofs of the same result.
-The first proof is a specialization of a proof of the Lindemann– Weierstrass theorem given by Beukers, Bézivin and Roba [2].
+We remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert.
+Then, using formal power series, we give two algebraic semiformal proofs of this result.
+The first proof is a specialization of the proof of the Lindemann–Weierstrass theorem found by Beukers, Bézivin and Robba [2].
 The second proof uses improper integrals of formal power series and is due to this author.
-At the beginning, we explain what “semiformal” means and indicate yet another analytical proof of the transcendence of $\mathrm{e}$.
-Like the two algebraic proofs, it avoids uncountable sets.
+We explain what “semiformal” means and indicate yet another analytical proof of the transcendence of $\mathrm{e}$.
+Like the two algebraic proofs, this proof avoids uncountable sets.
 It is due to this author in [7].
 1 Hilbert’s proof that $\mathrm{e}$ is transcendental. Why this proof is not completely satisfactory. How to fix it
-Hilbert’s proof [3] of the transcendence of the number $\mathrm{e}$ goes as follows.
+The transcendence of the number $\mathrm{e}$ was first proven by Hermite [4].
+Hilbert’s proof [5] goes as follows.
 We suppose for the contradiction that there exists $n+1$, $n\in\mathbb{N}$ ($\{1,2,\dots\}$), integers $a_{0}$, $a_{1}$, $\dots$, $a_{n}$, not all of them 0, such that
@@
 $p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}\,$,
-with $n$ as in $P$. We have the following expansions starting from the lowest powers of $x$:
-$p_{r}(x)=(-1)^{n(r+1)}(n!)^{r+1}\cdot x^{r}+\bullet\cdot x^{r+1}+\dots$
+with $n$ as in $P$. We have the expansions
+$p_{r}(x)=(-1)^{n(r+1)}(n!)^{r+1}\cdot x^{r}+ax^{r+1}+\dots$
 and
-$p_{r}(x+j)=\bullet\cdot x^{r+1}+\bullet\cdot x^{r+2}+\dots$
-for $j=1,2,\dots,n$ and some $\bullet\in\mathbb{Z}$ (the integers). We consider integrals
+$p_{r}(x+j)=bx^{r+1}+cx^{r+2}+\dots\,$,
+where $j=1,2,\dots,n$ and $a,b,c,\dots$ are in $\mathbb{Z}$ (the integers). We consider improper integrals
 $I_{r}=\int_{0}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x\,.$}
-Using additivity of the improper integral $\int_{0}^{+\infty}$ we express the product $P\cdot I_{r}$ as
-$0$ $=$ $P\cdot I_{r}=A_{r}+B_{r}$
+Using the additivity of improper integrals, we express the product $PI_{r}$ as
+$0$ $=$ $PI_{r}=A_{r}+B_{r}$
 $:=$ ${\textstyle\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x+\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x}\,.$
@@
 $|A_{r}|\leq c^{r}$
-where $c>1$ is a constant depending only on the coefficients $a_{0}$, $a_{1}$, $\dots$, $a_{n}$.
-The heart of the proof is the determination of the second quantity
+where $c\geq 1$ is a constant (depending only on the coefficients $a_{0}$, $a_{1}$, $\dots$, $a_{n}$).
+We determine the second quantity
 $B_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x\,.$}
-We claim that $B_{r}$ is an integral multiple of $r!$ and that $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. Since $c^{r}/r!\to 0$ as $r\to\infty$ for any constant $c>1$, the above equalities
+We claim that $B_{r}$ is an integral multiple of $r!$ and that $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$.
+Since $c^{r}/r!\to 0$ as $r\to\infty$ for any constant $c\geq 1$, the above equalities
 $A_{r}+B_{r}=0,\ r\in\mathbb{N}\,$,
-are contradictory and we are done. In order to prove the claim, we use the integral identity that
+are contradictory and we are done. We prove the claim by means of the integral identity that
 $\int_{0}^{+\infty}x^{n}\mathrm{e}^{-x}\,\mathrm{d}x=n!$}
 for every $n\in\mathbb{N}_{0}$ ($\{0,1,\dots\}$).
-It is due to Euler and is easy to establish by induction on $n$ via integration by parts.
-By the linearity of the improper integral $\int_{0}^{+\infty}$, for any integers $b_{0}$, $b_{1}$, $\dots$, $b_{m}$ with $m\in\mathbb{N}_{0}$ we have more generally that
+The identity is due to Euler and is easy to establish by induction on $n$ via integration by parts.
+Using the linearity of improper integrals, we obtain for any integers $b_{0}$, $b_{1}$, $\dots$, $b_{m}$ with $m\in\mathbb{N}_{0}$ the more general identity that
 $\int_{0}^{+\infty}\big(\sum_{j=0}^{m}b_{j}x^{j}\big)\mathrm{e}^{-x}\,\mathrm{d}x=\sum_{j=0}^{m}b_{j}\cdot j!\,.$}
-Then
+Thus
 $B_{r}$ $=$ ${\textstyle\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x}$
@@
 In the third line, we changed the variable to y=xjy=x-j.
-In the fourth line, we used the above generalization of Euler’s identity and the expansions of $p_{r}(x+j)$, $j=0,1,\dots,n$, in lowest powers of $x$.
+In the fourth line, we used the above generalization of Euler’s identity and the expansions of $p_{r}(x+j)$, $j=0,1,\dots,n$.
 It is clear that $B_{r}\in\mathbb{Z}$ and is divisible by $r!$.
@@
 $\Box$
-Interestingly, Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof that uses only proper integrals $\int_{a}^{b}$.
-Hilbert’s proof is elegant and short, but it is not completely satisfactory because it employs uncountable sets; such as, for example, the integrands in Euler’s identity.
-We explain how to fix it and how to prove the transcendence of the number $\mathrm{e}$ so that two conditions are satisfied.
+Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof using only proper integrals $\int_{a}^{b}$. A similar proof is given in Conrad [3].
+Hilbert’s proof is elegant and short, but it has the unsatisfactory feature that it employs uncountable sets, for example, the integrands in Euler’s identity.
+We explain how to improve upon it by proving the transcendence of $\mathrm{e}$ so that two conditions hold.
 Only hereditarily at most countable (HMC) sets are directly used in the proof.
 No technical notion of mathematical logic such as, for example, the notion of a formula, is used in the proof.
-A set $x$ is HMC if for every index $n\in\mathbb{N}_{0}$ and every chain of sets
+We call proofs satisfying these two conditions AB proofs. A set $x$ is HMC if for every index $n\in\mathbb{N}_{0}$ and every chain of sets
 $x_{n}\in x_{n-1}\in\dots\in x_{0}=x\,$,
-the set $x_{n}$ is finite or countable.
-As for uses of mathematical logic, we quote a referee of [4]: “That basic real analysis is essentially a countable endeavor has been well-known for over 100 years.
-My secondhand understanding is that Weyl articulated this very well in Das Kontinuum, which nowadays we would say developed analysis in $\mathrm{ACA}_{0}$.” The reader can check [5] for $\mathrm{ACA}_{0}$.
-Our more demanding goal is to obtain a proof of the transcendence of $\mathrm{e}$ satisfying besides (A) also (B), that is, such that no HMC set and no mathematical logic is used.
-There are (at least) three ways to achieve this goal.
+the set $x_{n}$ is at most countable (finite or countable).
+As for uses of mathematical logic in analysis, we quote a referee of [7]: “That basic real analysis is essentially a countable endeavor has been well-known for over 100 years.
+My secondhand understanding is that Weyl articulated this very well in Das Kontinuum, which nowadays we would say developed analysis in $\mathrm{ACA}_{0}$”.
+The reader can check [8] for $\mathrm{ACA}_{0}$.
+Our more ambitious and demanding goal is to obtain an AB proof of the transcendence of $\mathrm{e}$.
+In a related result, Kaibkhanov and Skopenkov [6] gave an AB proof of the transcendence of the Mahler number $\sum_{n\geq 0}2^{-2^{n}}$ based on binary expansions.
+There are at least three ways to such proof.
 The first way is to replace real functions of the type $f\colon[a,b]\to\mathbb{R}$, where $a<b$ are real numbers and $\mathbb{R}$ is the set of real numbers, with the functions
 $f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}\,$,
-where $\mathbb{Q}$ is the (HMC) set of fractions.
-If real numbers are represented by Dedekind cuts, any such function $f$ is an HMC set.
-We realized the first way in [4].
-Is there any difference in treatment of the two kinds of functions?
+where $\mathbb{Q}$ is the set of fractions (if real numbers are Dedekind cuts, every such function $f$ is an HMC set), and to develop a sufficient fragment of real analysis for such functions.
+We realized this way in [7].
+Is there any difference in the treatment of the two kinds of functions?
 The above mentioned referee thinks not: “Continuous functions on the reals are determined by their values on the rationals, so everything works out in the usual way.
-I don’t see how what the article is doing is so very different from developing analysis in (full) second-order arithmetic.” There is an obvious meta-argument refuting the referee’s first claim: if it were the case that “everything works out in the usual way,” then some enthusiasts would be teaching courses in analysis based on functions of the type $f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}$ for years.
+I don’t see how what the article is doing is so very different from developing analysis in (full) second-order arithmetic”.
+An obvious meta-argument refutes the first claim: if it were the case that “everything works out in the usual way”, then some enthusiasts would be teaching courses in analysis based on functions of the type $f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}$ for years.
 Since you have not heard about such courses, there must be some difficulties with this approach.
@@
 The interested reader is referred to [7] for details.
-We include the referee’s second claim just to reiterate that our approach in [4] satisfies condition (B).
-In [4], we do not use technical notions of mathematical logic at all; therefore, our approach is very different from any countable treatment of real analysis offered by mathematical logic, such as, to give besides Das Kontinuum more modern references, [7] or Weaver’s forthcoming book [8].
-We included this analytical digression to place the algebraic results of the present article in context.
-We present two algebraic proofs of the transcendence of $\mathrm{e}$ satisfying conditions (A) and (B).
-They are based on algebraic manipulations with formal power series.
+The second claim is refuted by the fact that in [7] we provide an AB proof.
+We do not use any technical notion of mathematical logic, and our approach in [7] is therefore very different from the countable treatments of real analysis offered by mathematical logic.
+For examples of the latter approaches, see Weaver [10, 11].
+This analytical digression places the algebraic results of the present article in context.
+These results are two AB proofs of the transcendence of $\mathrm{e}$.
+Both are based on formal power series.
 The first proof in Section 2 is a specialization of the proof in [2].
 The second proof in Section 3 is due to this author.
-We presented it in early versions of [4], but we rework it here.
-The word “directly” in condition (A) means that non-HMC sets may appear in the proof only by mention, as “harbors” of the main players.
-We may say, for example, that for every two real sequences $(a_{n}),(b_{n})\subset\mathbb{R}$ we have $\dots$; it is clear that $\mathbb{R}$ is used only indirectly and can be easily formulated away.
-We allow such uses of non-HMC sets for the sake of natural and easily understandable formulations.
-In conclusion of this introduction, we review formal power series and explain the meaning of the adjective “semiformal”.
-If $R$ is an (integral) domain, we write $R[[x]]$ for the (usually non-HMC) set of maps $f\colon\mathbb{N}_{0}\to R$.
-They are traditionally written as infinite formal sums
-$f=f(x)=\sum_{n\geq 0}a_{n}x^{n}$}\,,
-where $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also by $x^{n}]f$. The set $R[[x]]$ is endowed with the operations
-${\textstyle\sum_{n\geq 0}a_{n}x^{n}+\sum_{n\geq 0}a_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}(a_{n}+b_{n})x^{n}\,\text{ and}}$
-${\textstyle\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}a_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}\,,}$
+We presented it already in early versions of [7], and we rework it here.
+The word “directly” in condition (A) means that non-HMC sets may appear in the proof only as supersets of the main, directly used sets, which have to be HMC.
+We may say, for example, that for every two sequences $(a_{n}),(b_{n})\subset\mathbb{R}$ we have $\dots\;$.
+It is clear that the set $\mathbb{R}$ is mentioned only for convenience and can be easily formulated away.
+We allow such uses of non-HMC sets in AB proofs for the sake of natural and easily legible formulations.
+In conclusion of the introduction, we review formal power series and explain the meaning of the adjective “semiformal”.
+If $R$ is an (integral) domain, we write $R[[x]]$ for the uncountable set of maps $f\colon\mathbb{N}_{0}\to R$.
+They are written as infinite formal sums
+$f=f(x)=\sum_{n\geq 0}a_{n}x^{n}$}
+with $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also as $x^{n}]f$. The set $R[[x]]$ is endowed with the operations
+${\textstyle f(x)+g(x)=\sum_{n\geq 0}a_{n}x^{n}+\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}(a_{n}+b_{n})x^{n}\,\text{ and}}$
+${\textstyle f(x)\cdot g(x)=\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}\,,}$
 and neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. Then
 R[[x]]=R[[x]], $\langle R[[x]],\,0,\,1$$,\cdot\rangle$
-is a domain and $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is (formal) derivative,
+is a domain and $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is the (formal) derivative,
 $\big(\sum_{n\geq 0}a_{n}x^{n}\big)^{\prime}=\sum_{n\geq 0}(n+1)_{R}\cdot a_{n+1}x^{n}\,.$}
-So for derivatives the characteristic of $R$ becomes relevant. Less often one uses the (formal) primitive
-$\int\sum_{n\geq 0}a_{n}x^{n}=\sum_{n\geq 0}\frac{a_{n}}{n+1}x^{n+1}=\sum_{n\geq 1}\frac{a_{n-1}}{n}x^{n}\,$,}
-but in this article, it is more important than the derivative.
-If $K$ is the normed field $K=\mathbb{R}$ or $K=\mathbb{C}$ (complex numbers), we denote by $K\{x\}$ the subdomain of $K[[x]]$ formed by convergent formal power series $f(x)=\sum_{n\geq 0}a_{n}x_{n}$ that satisfy
-$\lim_{n\to\infty}a_{n}c^{n}=0$
-for every $c\in K$. In other words, $f(x)$ has an infinite radius of convergence.
-The binary and unary operations on $R[[x]]$ that we have listed so far, ++, $\cdot$, $(\dots)^{\prime}$, and $\int(\dots)$, are formal in the sense that, for example, $x_{n}](f\cdot g)$ is determined by finitely many coefficients in $f$ and $g$, and similarly for the other operations.
-In formal arguments involving elements $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$ it is also forbidden to use values
-$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k}$}
-of $f(x)$ on elements $c\in K$.
-It is still an open problem to find a fully formal proof of the transcendence of $\mathrm{e}$ based on formal power series; or at least no such proof is known to this author.
-However, in the next two sections we present two such proofs if formality is relaxed to quasiformality.
-An argument involving elements of $K[[x]]$ is quasiformal if it uses only formal and quasiformal operations, and for every $f(x)\in K[[x]]$ appearing in it only at most countably many values $f(c)$ for $c\in K$ are used.
-Quasiformal operations allow uses of limit transitions in $K$ and of at most countably many values $f(c)$.
-Quasiformal operations on $\mathbb{R}[[x]]$ are introduced, for example, in Definitions 3.3 and 3.11.
+Less often one uses the (formal) primitive
+$\int\sum_{n\geq 0}a_{n}x^{n}=\sum_{n\geq 0}\frac{a_{n}}{n+1}x^{n+1}=\sum_{n\geq 1}\frac{a_{n-1}}{n}x^{n}\,.$}
+In this article, $\int$ is more important than the derivative.
+If $K$ is the normed field $K=\mathbb{R}$ or $K=\mathbb{C}$ (complex numbers), we denote by $K\{x\}$ the subdomain of $K[[x]]$ formed by the convergent formal power series $f(x)=\sum_{n\geq 0}a_{n}x_{n}$.
+They satisfy for every $c\in K$ that
+$\lim_{n\to\infty}a_{n}c^{n}=0_{K}\,.$
+In other words, $f(x)\in K\{x\}$ iff $f(x)$ has an infinite radius of convergence.
+The binary and unary operations on $R[[x]]$ listed so far, ++, $\cdot$, $(\cdot)^{\prime}$, and $\int(\cdot)$, are formal in the sense that, for example, for every $n\in\mathbb{N}_{0}$ the coefficient
+$x^{n}](f\cdot g)$
+is determined by only finitely many coefficients in $f$ and $g$, and the same holds for the other operations.
+In formal proofs involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$, only formal operations may be used, and it is forbidden to use the functional values
+$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k},\ c\in K\,.$}
+It is an open problem to give a completely formal AB proof of the transcendence of $\mathrm{e}$ based on formal power series. No such proof is known to this author.
+However, we can present two such proofs in the next two sections if formality is relaxed to semiformality.
+Semiformal operations (in Section 3 we introduce several of them) are defined on $K\{x\}$, may use limit transitions, and may use at most countably many values $f(c)$ for $f\in K[[x]]$ and $c\in K$.
+Semiformal proofs concerning elements in $K[[x]]$ are proofs that satisfy the following condition.
+Only formal and semiformal operations on $K[[x]]$ are used. For every formal power series involved in the proof, only at most countably many of its values are used.
+In Sections 2 and 3 we give two AB &\& SF proofs of the transcendence of $\mathrm{e}$.
 2 Specializing the proof of Beukers, Bézivin and Robba
 The Lindemann–Weierstrass theorem says that
-$b_{1}\mathrm{e}^{a_{1}}+b_{2}\mathrm{e}^{a_{2}}+\dots+b_{t}\mathrm{e}^{a_{t}}\neq 0$
+$b_{1}\mathrm{e}^{\alpha_{1}}+b_{2}\mathrm{e}^{\alpha_{2}}+\dots+b_{t}\mathrm{e}^{\alpha_{t}}\neq 0$
 whenever $b_{j}$ and $\alpha_{j}$ are $2t$, with $t\in\mathbb{N}$, algebraic numbers in $\mathbb{C}$ such that every $b_{j}\neq 0$ and the $\alpha_{j}$ are mutually distinct.
@@
 6–8] for an analytic proof.
-Using formal power series, Beukers, Bézivin and Robba gave in [2] another proof of the LW theorem.
+Using formal power series, Beukers, Bézivin, and Robba [2] obtained another proof of the LW theorem.
 In this section, we specialize their proof to $b_{j},\alpha_{j}\in\mathbb{Z}$, which gives the transcendence of $\mathrm{e}$.
-In order to clearly describe logical structure of the proof, we completely reorganize the exposition in [2].
-At the end of this section, we explain why this proof is semiformal (and not formal), and why it satisfies conditions (A) and (B).
-We begin by recalling that a formal power series $f(x)$ in $\mathbb{C}[[x]]$ is rational if there exist two polynomials $p(x),q(x)\in\mathbb{C}[x]$ such that $q(0)=1$ and $q(x)f(x)=p(x)$; we may assume that $p(x)$ and $q(x)$ are coprime.
+In order to clearly present the logical structure of their proof, we reorganize the exposition in [2].
+We first recall that a formal power series $f(x)$ in $\mathbb{C}[[x]]$ is rational if there exist two polynomials $p(x),q(x)\in\mathbb{C}[x]$ such that $q(0)=1$ and $q(x)f(x)=p(x)$; we may assume that $p(x)$ and $q(x)$ are coprime.
 If $f(x)$ and $g(x)$ in $\mathbb{C}[[x]]$ are rational, then so are $f(x)+g(x)$, $f(x)g(x)$, and $f^{\prime}(x)$.
-Any rational $f(x)\in\mathbb{C}[[x]]$ has a unique decomposition in partial fractions
+Any rational formal power series $f(x)\in\mathbb{C}[[x]]$ has a unique decomposition into partial fractions
 $f(x)=r(x)+\sum_{j=1}^{t}\sum_{i=1}^{m_{j}}\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}\,$,}
-where $r(x)\in\mathbb{C}[x]$, $t\in\mathbb{N}_{0}$, $m_{j}\in\mathbb{N}$, $\beta_{j,i},\alpha_{j}\in\mathbb{C}$ with $\beta_{j,m_{j}}\neq 0$ and the $\alpha_{j}$ are nonzero and mutually distinct; for $t=0$, the double sum is interpreted as 0, and the numbers $m_{j}$ and $\alpha_{j}$ come from the factorization
+where $r(x)\in\mathbb{C}[x]$, $t\in\mathbb{N}_{0}$, $m_{j}\in\mathbb{N}$, $\beta_{j,i},\alpha_{j}\in\mathbb{C}$ with $\beta_{j,m_{j}}\neq 0$, and the $\alpha_{j}$ are nonzero and mutually distinct.
+For $t=0$, the double sum is interpreted as 0, and the numbers $m_{j}$ and $\alpha_{j}$ come from the factorization
 $q(x)=\prod_{j=1}^{t}(1-\alpha_{j}x)^{m_{j}}\,.$}
 The partial fractions are the formal power series
-$\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}=\beta_{j,i}\sum_{n\geq 0}\binom{-i}{n}\alpha_{j}^{n}x^{n}\,.$}
+$\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}=\sum_{n\geq 0}\binom{-i}{n}\beta_{j,i}\alpha_{j}^{n}\cdot x^{n}\,.$}
 We say that $1/\alpha_{j}$ is a pole of $f(x)$ with order $m_{j}$.
@@
 Proof.
-This follows from the definition of the decomposition of a rational formal power series in partial fractions.
-The product $p(x)A(x)$ possibly only decreases the orders of the poles of $A(x)$, or cancels them and does not create any new poles.
-The product $cx^{m}A^{\prime}(x)$ preserves all poles of $A(x)$, increases their orders by $1$, and does not create any new poles.
+The product $p(x)A(x)$ only possibly decreases the orders of the poles of $A(x)$ or cancels them and does not create any new poles.
+The product and derivative $cx^{m}A^{\prime}(x)$ preserve all poles of $A(x)$, increase their orders by $1$, and do not create any new poles.
 The sum $p(x)A(x)+cx^{m}A^{\prime}(x)$ merges the two sets of poles.
-Therefore, the resulting set of poles with their orders is that of $cx^{m}A^{\prime}(x)$.
+It follows that the resulting set of poles with their orders is that of $cx^{m}A^{\prime}(x)$.
 $\Box$
-We prove that $\mathrm{e}$ is transcendental by means of the following theorem on rationality of formal power series, which is extracted from the arguments in [2].
+We extracted the following theorem on rational formal power series from the arguments in [2].
 Theorem 2.2
@@
 $v_{n}(k+1)=v_{n}(k)-a_{1}v_{n-1}(k)-\dots-a_{t}v_{n-t}(k)\,.$
-We establish for $n,k\in\mathbb{N}_{0}$ with $n\geq tk$ two properties of the numbers $v_{n}(k)$.
+We establish for every $n,k\in\mathbb{N}_{0}$ with $n\geq tk$ two properties of the numbers $v_{n}(k)$.
   1. 1.
-    We have bound $|v_{n}(k)|\leq cA^{n}C^{k}$ where $C=1+|a_{1}|+\dots+|a_{t}|$ and $c>0$ is a constant.
+    We have the bound $|v_{n}(k)|\leq cA^{n}C^{k}$ where $C=1+|a_{1}|+\dots+|a_{t}|$ and $c\geq 0$ is a constant.
   2. 2.
     $k!\,|\,v_{n}(k)$.
-The bound in item 1 follows by induction on $k$ and the assumed bound on $v_{n}=v_{n}(0)$.
-For $k=0$ the bound clearly holds.
+The bound in item 1 follows by induction on $k$ from the assumed bound on $v_{n}=v_{n}(0)$.
+For $k=0$ the bound holds because it is the assumed bound on $v_{n}$.
 Let $n,k\in\mathbb{N}_{0}$ with $n\geq t(k+1)$.
@@
 $v_{n}(k)=v_{n}(k+1)+a_{1}v_{n-1}(k)+\dots+a_{t}v_{n-t}(k),\ n\geq t\,$,
-of the above recurrence, we show that for a fixed $k\in\mathbb{N}$ we have $v_{n}(k)=0$ for every large $n$.
+of the above recurrence, we show that there exists a $k\in\mathbb{N}$ such that $v_{n}(k)=0$ for every large $n$.
 This will prove that the formal power series $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ is rational.
@@
     For every $i=0,1,\dots,t$, the norm $\|\beta_{i}\|<\|\gamma\|$.
-From these three properties and from the modified recurrence, it follows by induction on the norm that $v_{n}(k)=0$ for every $\langle k,n\rangle\in M$.
-This concludes the proof because we see that every $k\geq k_{0}$ has the required property.
+From these three properties, from the property of the set $N$ and from the modified recurrence, it follows by induction on the norm that $v_{n}(k)=0$ for every $\langle k,n\rangle\in M$.
+This concludes the proof because we see that every number $k\geq k_{0}$ has the required property.
 $\Box$
@@
 $\leq$ ${\textstyle cA^{n}\sum_{j\geq 0}\frac{A^{j}}{j!}=c\,\mathrm{e}^{A}\cdot A^{n}\,,}$
-where $c>0$ is an absolute constant.
+where $c\geq 0$ is an absolute constant.
 Thus $v_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$.
@@
 contradicts Proposition 2.1 because by Theorem 2.2 the formal power series $V(x)$ is rational. $\Box$
-We reiterate that the proof in [2] is more general as it establishes the full LW theorem.
-In this situation, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of.
-The previous proof works with formal power series, but it is only semiformal and not formal because the key player is the infinite sum
+We repeat that the proof in [2] is more general as it establishes the full LW theorem.
+In that proof, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of.
+The previous proof works with formal power series, but it is semiformal and not formal because the key player is the infinite sum
 $\sum_{n\geq 0}\frac{u_{n}}{n!}=0$}
-which is the value $f(1)$ of the formal power series $f(x)=\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}$. In fact, this is the only non-formal feature of it.
+which is the value $f(1)$ of the formal power series $f(x)=\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}$. In fact, this is the only non-formal component in an otherwise formal proof.
 It is clear that the previous proof satisfies condition (A), which states that all directly used sets are HMC.
 It works directly only with complex formal power series, which are maps from $\mathbb{N}_{0}$ to $\mathbb{C}$.
-As explained in Section 1, non-HMC sets like $\mathbb{C}[[x]]$ are not directly used in the proof.
-Condition (B), which states that no formalism of mathematical logic is used, is clearly satisfied as well.
-Although the article [2] of Beukers, Bézivin, and Robba is a pioneering one, as it is the first proof of the transcendence of the number $\mathrm{e}$ (in fact, of course, of the full LW theorem) that satisfies conditions (A) and (B), it is little known; in the database of Mathematical Reviews, it currently has only 11 citation (January 2026).
+As explained in Section 1, non-HMC sets like $\mathbb{C}[[x]]$ are not really needed in the proof.
+Condition (B), which states that no formalism of mathematical logic is used, is satisfied as well.
+Although the article [2] of Beukers, Bézivin, and Robba is a pioneering one, as it is the first AB proof of the transcendence of the number $\mathrm{e}$ (in fact, of course, of the full LW theorem), it is little known; in the database of Mathematical Reviews, it currently has only 11 citation (January 2026).
 3 The proof by formal integration
-We present the second semiformal proof of the transcendence of $\mathrm{e}$.
+We present the second AB &\& SF proof of the transcendence of $\mathrm{e}$.
 It is due to this author and first appeared in early versions of [7].
-The idea is simply to translate Hilbert’s proof, which we described at the beginning, into the language of formal power series.
-At the end of the section we explain why the proof is semiformal and why conditions (A) and (B) are satisfied.
+The idea is to translate Hilbert’s proof, which we described at the beginning, into the language of formal power series.
 We work in the domains of formal power series $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$.
-We introduce additional operations on $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$ besides those we reviewed in Section 1.
+We introduce further operations on $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$ besides those we mentioned in Section 1.
 For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
 $f(ax)=\sum_{n\geq 0}a^{n}a_{n}x^{n}\,.$}
-This is a formal operation. Clearly, for every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
+The operation $f(x)\mapsto f(ax)$, which we call the constant inner product, is formal. Clearly, for every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
 $(af+bg)(cx)=af(cx)+bg(cx)\,\text{ and }\,(fg)(cx)=f(cx)g(cx)\,.$
-If $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ then $f(0x)=a_{0}$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$.
-In Section 1 we defined the derivative $f^{\prime}$ and the primitive $\int f$ of a formal power series $f\in\mathbb{R}[[x]]$.
-In the next proposition, whose straightforward proof we omit, we collect properties of these operations.
+If $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ then $f(0x)=a_{0}$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
+In the next proposition, whose straightforward proof we omit, we collect properties of the operations $f\mapsto f^{\prime}$ and $f\mapsto\int f$.
 Proposition 3.1
@@
 $f(x+b)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)x^{n}\;.$}
-The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent, and $f(x+b)\in\mathbb{R}\{x\}$.
-It is easy to see that $f(x+0)=f(x)$ and that if $f(x)\in\mathbb{R}[x]$ then $f(x+b)$ is the usual substitution in a polynomial.
-In the next two propositions we show that semiformal shifts behave in the expected way.
+The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent.
+It is easy to show that $f(x+b)\in\mathbb{R}\{x\}$.
+Clearly, $f(x+0)=f(x)$ and if $f(x)\in\mathbb{R}[x]$ then $f(x)\mapsto f(x+b)$ is the usual polynomial substitution.
+We combine the operations of constant inner product and semiformal shift by means of brackets.
+For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
+In the next two propositions, we show that semiformal shifts behave in desirable ways.
 Proposition 3.4
-Let $f(x)\in\mathbb{R}\{x\}$, $c,d\in\mathbb{R}$ and $g(x)=f(x+c)$. Then
-g(x+d)=f(x+(c+d)).g(x+d)=f(x+(c+d))\;.
-Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $c$, $d$ and $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ be as stated. Then
-$b_{n}=\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,g(x+d)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}\big)x^{n}\,.$}
-Thus the coefficient of $x^{n}$ in $g(x+d)$ equals
-${\textstyle\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}=\sum_{m\geq n}\binom{m}{n}\big(\sum_{l\geq m}\binom{l}{m}a_{l}c^{l-m}\big)d^{m-n}=}$
+Let $f(x)\in\mathbb{R}\{x\}$ and $c,d\in\mathbb{R}$. Then
+f((x+c)+d)=f(x+(c+d)).f((x+c)+d)=f(x+(c+d))\,.
+Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ and $c,d$ be as stated, and let $f(x+c)=\sum_{n\geq 0}b_{n}x^{n}$. Then
+$b_{n}=\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,f((x+c)+d)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}\big)x^{n}\,.$}
+Thus
+${\textstyle[x^{n}]f((x+c)+d)=\sum_{m\geq n}\binom{m}{n}\big(\sum_{l\geq m}\binom{l}{m}a_{l}c^{l-m}\big)d^{m-n}=}$
 $={\textstyle\sum_{m\geq n}\sum_{l\geq m}\frac{l!}{n!(m-n)!(l-m)!}a_{l}c^{l-m}d^{m-n}=}$
 ${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{m=n}^{l}\binom{l-n}{l-m}c^{l-m}d^{m-n}=}$
-${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{j=0}^{l-n}\binom{l-n}{j}c^{j}d^{l-n-j}=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}\,,}$
-which is the coefficient of $x^{n}$ in $f(x+(c+d))$.
-In the third line, we changed the order of summation.
-This is allowed because $f(x)$ is convergent.
-In the fourth line, we used the binomial theorem.
-$\Box$
+${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{j=0}^{l-n}\binom{l-n}{j}c^{j}d^{l-n-j}=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}=}$
+$=[x^{n}]f(x+(c+d))\,.$
+In the third line, we changed the order of summation. This is allowed because $f(x)$ is convergent. In the fourth line, we used the binomial theorem. $\Box$
 Proposition 3.5
-Let $f,g\in\mathbb{R}\{x\}$ and $c\in\mathbb{R}$. Then
+Let $f(x),g(x)\in\mathbb{R}\{x\}$ and $c\in\mathbb{R}$. Then
 $(f\cdot g)(x+c)=f(x+c)\cdot g(x+c)\;.$
@@
 The coefficient of $x^{n}$ in the formal power series $f(x+c)$, respectively $g(x+c)$, equals $\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}$, respectively $\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}$.
-Thus the coefficient of $x^{n}$ in the product $f(x+c)\cdot g(x+c)$ equals
-${\textstyle\sum_{k=0}^{n}\big(\sum_{m\geq k}\binom{m}{k}a_{m}c^{m-k}\cdot\sum_{l\geq n-k}\binom{l}{n-k}b_{l}c^{l+k-n}\big)=}$
+Thus
+${\textstyle[x^{n}]f(x+c)g(x+c)=\sum_{k=0}^{n}\big(\sum_{m\geq k}\binom{m}{k}a_{m}c^{m-k}\sum_{l\geq n-k}\binom{l}{n-k}b_{l}c^{l+k-n}\big)}$
 ${\textstyle=\sum_{k=0}^{n}\sum_{m\geq k}\sum_{l\geq n-k}\binom{m}{k}\binom{l}{n-k}a_{m}b_{l}c^{l+m-n}=}$
 ${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\sum_{k=0}^{n}\binom{m}{k}\binom{l}{n-k}a_{m}b_{l}c^{l+m-n}=}$
-${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\binom{m+l}{n}a_{m}b_{l}c^{l+m-n}=\sum_{j\geq n}\binom{j}{n}\big(\sum_{m=0}^{j}a_{m}b_{j-m}\big)c^{j-n}\,,}$
-which is the coefficient of $x^{n}$ in $(f\cdot g)(x+c)$.
-In the third line, we changed the order of summation.
-This is allowed since $f(x)$ is convergent.
-In the fourth line, we used Vandermonde’s identity ([6]).
-$\Box$
-We prove a semiformal exponential identity.
-Theorem 3.6
-Let $a,b\in\mathbb{R}$ with $a\neq 0$ and $f(x)=\mathrm{Exp}\big(x+\frac{b}{a}\big)$. Then
-$f(ax)=\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$
-Proof. Let $a$, $b$ and $f(x)$ be as stated. Then, indeed,
-$f(ax)$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}(a^{m}/m!)(b/a)^{m-n}\big)x^{n}}$
-$=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}=\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)\frac{a^{n}}{n!}x^{n}}$
+${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\binom{m+l}{n}a_{m}b_{l}c^{l+m-n}=\sum_{j\geq n}\binom{j}{n}\big(\sum_{m=0}^{j}a_{m}b_{j-m}\big)c^{j-n}=}$
+$=[x^{n}](fg)(x+c)\,.$
+In the third line, we changed the order of summation. This is allowed since $f(x)$ is convergent. In the fourth line, we used Vandermonde’s identity ([9]). $\Box$
+The following semiformal exponential identity is somewhat counter-intuitive (at first we got it wrong).
+Proposition 3.6
+Let $a,b\in\mathbb{R}$ with $a\neq 0$. Then
+$\mathrm{Exp}\big((ax)+\frac{b}{a}\big)=\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$}
+Proof. Let $a$, $b$ be as stated, and let $f(x)=\mathrm{Exp}(ax)$. Then
+${\textstyle f\big(x+\frac{b}{a}\big)}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}(a^{m}/m!)(b/a)^{m-n}\big)x^{n}}$
+$=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}=\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)(a^{n}/n!)x^{n}}$
 $=$ $\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$
 $\Box$
-We introduce the Newton integral of convergent formal power series.
+We will use this identity only with a=1a=-1.
+We introduce Newton integrals of convergent formal power series.
 Definition 3.7 (Newton integral of fps)
@@
 of the values of the formal primitive of $f$ at the numbers $v$ and $u$.
-The integral is always defined by item 3 of Proposition 3.1 because $\int f\in\mathbb{R}\{x\}$.
+By item 3 of Proposition 3.1, $\int f\in\mathbb{R}\{x\}$ and the integral is always defined.
 The Newton integral is a semiformal operation because it uses two values of a formal power series.
-We establish for it three integral transformations: linear combination, additive combination, and shift of the integration interval.
-The first two are quite straightforward.
+We prove three integral transformations for it: linear combination, additive decomposition, and shift of the integration interval.
+The first two transformations are straightforward.
 Proposition 3.8
@@
 This is allowed since the involved series absolutely converge.
-In the fourth line, we add and subtract the term $b^{m+1}$ in the two inner sums.
-In the fifth line, we use the binomial theorem.
-$\Box$
-We introduce the improper Newton integral of formal power series.
+In the fourth line, we add
+$0=\sum_{m\geq 0}\frac{a_{m}}{m+1}\binom{m+1}{0}b^{m+1}v^{0}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\binom{m+1}{0}b^{m+1}u^{0}\,.$}
+In the fifth line, we use the binomial theorem. $\Box$
+We introduce improper Newton integrals of convergent formal power series.
 Definition 3.11 (improper Newton $\int$ of fps)
-Let $u\in\mathbb{R}$ and $f(x)\in\mathbb{R}\{x\}$. If for every real sequence $(b_{n})$ going to $\infty$ the limit
+Let $u\in\mathbb{R}$ and $f(x)\in\mathbb{R}\{x\}$. If for every real sequence $(b_{n})$ going to $\infty$ the proper limit
 $L=\lim_{n\to\infty}\int_{u}^{b_{n}}f\ \ (\in\mathbb{R})$}
@@
 $\int_{u}^{+\infty}f=\int_{u}^{+\infty}f(x):=L\,.$}
-We call this integral the improper Newton integral from $u$ to $\infty$ of the formal power series $f(x)$.
+We call $L$ the improper Newton integral of $f(x)$ from $u$ to $\infty$.
 We leave the proof of the independence of the limit $L$ on the sequences $(b_{n})$ to the reader as an exercise.
-We obtain for the improper integral the same three integral transformations as before.
+We prove the same three integral transformations as before.
 Proposition 3.12
@@
 But then
-${\textstyle\int_{u}^{+\infty}f=}$
-${\textstyle=L-\big(\int f\big)(u)=\big(\int f\big)(v)-\big(\int f\big)(u)+L-\big(\int f\big)(v)=}$
+${\textstyle\int_{u}^{+\infty}f=L-\big(\int f\big)(u)=\big(\int f\big)(v)-\big(\int f\big)(u)+L-\big(\int f\big)(v)=}$
 ${\textstyle=\int_{u}^{v}f+\int_{v}^{+\infty}f\,.}$
 With the assumption of existence of $\int_{v}^{+\infty}f$ we argue similarly. $\Box$
-The third transformation is again less immediate.
+The third transformation is immediate as well.
 Proposition 3.14
-Let $f\in\mathbb{R}\{x\}$ and $u,b\in\mathbb{R}$. The equality
+Let $f(x)\in\mathbb{R}\{x\}$ and $u,b\in\mathbb{R}$. The equality
 $\int_{u}^{+\infty}f(x+b)=\int_{u+b}^{+\infty}f(x)$}
@@
 Proof.
-Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $u$ and $b$ be as stated.
-Let $(c_{n})$ be any real sequence converging to $\infty$, and let $L$ and $R$ be the integrals on the left-hand side and the right-hand side, respectively.
-It suffices to show the implication $\exists\,R\Rightarrow R=L$; the opposite implication then follows.
-Indeed, we suppose that integral $L$ exists and write it as
-$L=\int_{u+b-b}^{+\infty}f(x+b)\,.$}
-We set $g(x)=f(x+b)$, use the assumed implication with b-b and get that
+Let $f(x)$ and $u,b$ be as stated.
+Let $(c_{n})$ be any real sequence going to $\infty$.
+We assume that the integral $I$ on the left-hand side exists, and prove that the integral on the right-hand side exists too and equals $I$.
+Indeed,
 […]
-Thus, the implication $\exists\,L\Rightarrow L=R$ holds as well.
-We suppose that integral $R$ exists. We replace the test sequence $(c_{n})$ with $(c_{n}+b)$ and get
-$R$ $=$ $\lim_{m\to\infty}{\textstyle\big(\sum_{n\geq 0}a_{n}\frac{(c_{m}+b)^{n+1}}{n+1}\big)-\sum_{n\geq 0}a_{n}\frac{(u+b)^{n+1}}{n+1}}$
-$=$ $\lim_{m\to\infty}{\textstyle\big(\sum_{l\geq 0}\sum_{n\geq l}\binom{n+1}{l+1}a_{n}b^{n-l}\frac{c_{m}^{l+1}}{n+1}\big)-\sum_{l\geq 0}\sum_{n\geq l}\binom{n+1}{l+1}a_{n}b^{n-l}\frac{u^{l+1}}{n+1}}$
-$=$ $\lim_{m\to\infty}{\textstyle\big(\sum_{l\geq 0}\sum_{n\geq l}\binom{n}{l}a_{n}b^{n-l}\frac{c_{m}^{l+1}}{l+1}\big)-\sum_{l\geq 0}\sum_{n\geq l}\binom{n}{l}a_{n}b^{n-l}\frac{u^{l+1}}{l+1}}$
-$=$ $\lim_{m\to\infty}{\textstyle\big(\int f(x+b)\big)(c_{m})}-{\textstyle\big(\int f(x+b)\big)(u)}=L\,.$
-In the second line, we used the binomial theorem, and we changed the order of summation, which is allowed by the convergence of $f(x)$.
-Also, we canceled the summands corresponding to $l+1=0$.
-In the third line, we used the transformation $\binom{l+1}{n+1}=\frac{l+1}{n+1}\binom{l}{n}$.
-In the fourth line, we use the definition of formal primitive and Definition 3.3.
+Similarly, if the integral on the right-hand side exists, then
+[…]
 $\Box$
-We need bounds on integrals
-$\int_{0}^{i}p(x)\mathrm{Exp}(-x)$}
-where $i\in\mathbb{N}_{0}$, $p(x)$ is an integral polynomial and $\mathrm{Exp}(-x)=\mathrm{Exp}((-1)x)$.
-In the semiformal approach, this appears problematic because one deduces the standard LM bound by considering all values of the integrand on $[0,i]$ or $0,i]\cap\mathbb{Q}$.
-Instead, we use the next substitute LM bound, which can be proven by means of just two values of the primitive.
+We need to bound Newton integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)$ is in $\mathbb{Z}[x]$. We easily obtain the next bound based on two values of the primitive.
 Proposition 3.15
-For every $i,k\in\mathbb{N}_{0}$,
+For every numbers $i,k\in\mathbb{N}_{0}$,
 $\big|\int_{0}^{i}x^{k}\cdot\mathrm{Exp}(-x)\big|\leq i^{k+1}\mathrm{e}^{i}\,.$}
 Proof. The integrand
-$\text{$\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!}x^{n+k}$ has the primitive $\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!(n+k+1)}x^{n+k+1}$}\,.$}
-Thus the Newton integral is in absolute value at most $\sum_{n=0}^{\infty}\frac{1}{n!}i^{n+k+1}+0=i^{k+1}\mathrm{e}^{i}$. $\Box$
-It is important that the bound is exponential and not factorial in $k$.
-The key step in Hilbert’s proof of the transcendence of $\mathrm{e}$ was Euler’s identity. Now we prove it in the setting of the improper Newton integral of formal power series.
-Theorem 3.16
+$\text{$\sum_{n\geq 0}\frac{(-1)^{n}}{n!}x^{n+k}$ has the primitive $\sum_{n\geq 0}\frac{(-1)^{n}}{n!(n+k+1)}x^{n+k+1}$}\,.$}
+Thus
+$\big|\int_{0}^{i}x^{k}\cdot\mathrm{Exp}(-x)\big|\leq\sum_{n\geq 0}\frac{1}{n!}i^{n+k+1}+0=i^{k+1}\mathrm{e}^{i}\,.$}
+$\Box$
+It is crucial that the bound is exponential and not factorial in $k$.
+The key step in Hilbert’s proof of the transcendence of $\mathrm{e}$ is Euler’s identity. We prove it in the language of improper Newton integrals of convergent formal power series.
+Proposition 3.16
 For every $k\in\mathbb{N}_{0}$, the integral
-$I_{k}=\int_{0}^{+\infty}x^{k}\cdot\mathrm{Exp}(-x)=\int_{0}^{+\infty}\sum_{n=k}^{\infty}\frac{(-1)^{n-k}}{(n-k)!}x^{n}=k!\,.$}
-Proof. We proceed by induction on $k$. Let $(a_{n})$ be any real sequence going to $\infty$. For $k=0$ we use that
-$\int\mathrm{Exp}(-x)=\sum_{n\geq 0}\frac{(-1)^{n}}{(n+1)!}x^{n+1}=1-\mathrm{Exp}(-x)\,.$}
-Hence
+$I_{k}=\int_{0}^{+\infty}x^{k}\cdot\mathrm{Exp}(-x)=\int_{0}^{+\infty}\sum_{n\geq k}\frac{(-1)^{n-k}}{(n-k)!}x^{n}=k!\,.$}
+Proof. We proceed by induction on $k$. Let $(a_{n})$ be any real sequence going to $\infty$. Since
+$\int\mathrm{Exp}(-x)=\sum_{n\geq 0}\frac{(-1)^{n}}{(n+1)!}x^{n+1}=1-\mathrm{Exp}(-x)\,$,}
+we have
 $I_{0}=\lim_{n\to\infty}{\textstyle(1-\mathrm{Exp}(-a_{n}))-(1-\mathrm{Exp}(-0))=1-0=1\,.}$
-For $k\in\mathbb{N}_{0}$, let $F_{k}=x^{k}\cdot\mathrm{Exp}(-x)$. Let $k\in\mathbb{N}$. By item 1 of Proposition 3.1 and the derivative of $\mathrm{Exp}(-x)$,
-$F_{k}^{\prime}=\big(x^{k}\cdot\mathrm{Exp}(-x)\big)^{\prime}=kx^{k-1}\cdot\mathrm{Exp}(-x)-x^{k}\cdot\mathrm{Exp}(-x)=k\cdot F_{k-1}-F_{k}\,.$
-By item 4 of Proposition 3.1, $\int F_{k}^{\prime}=F_{k}$. Thus by Proposition 3.12 and induction,
+Let $F_{k}=x^{k}\cdot\mathrm{Exp}(-x)$, $k\in\mathbb{N}_{0}$, and let $k\in\mathbb{N}$. By item 1 of Proposition 3.1 and the derivative of $\mathrm{Exp}(-x)$,
+$F_{k}^{\prime}=\big(x^{k}\cdot\mathrm{Exp}(-x)\big)^{\prime}=kx^{k-1}\cdot\mathrm{Exp}(-x)-x^{k}\cdot\mathrm{Exp}(-x)=k\cdot F_{k-1}-F_{k}$
+and $F_{k}=k\cdot F_{k-1}-F_{k}^{\prime}$. By item 4 of Proposition 3.1, $\int F_{k}^{\prime}=F_{k}-[x^{0}]F_{k}=F_{k}$. By Proposition 3.12 and induction,
 $I_{k}$ $=$ $\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k}}=k\cdot\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k-1}-}\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k}^{\prime}}$
-$=$ $k\cdot I_{k-1}-\lim_{n\to\infty}F_{k}(a_{n})+F_{k}(0)$
-$=$ $k\cdot(k-1)!-0+0=k!\;.$
+$=$ $k\cdot I_{k-1}-\lim_{n\to\infty}\big(F_{k}(a_{n})-F_{k}(0)\big)$
+$=$ $k\cdot(k-1)!-0=k!\;.$
 $\Box$
@@
 $a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0$
-where $a_{i}\in\mathbb{Z}$, $n\in\mathbb{N}$, and $a_{0}\neq 0$. We multiply this equality by the improper Newton integral (of formal power series)
+where $a_{i}\in\mathbb{Z}$, $n\in\mathbb{N}$, and $a_{0}\neq 0$. We multiply this equality by the improper Newton integral
 $\int_{0}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)\,$,}
-where $p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$ (as in Section 1). This integral exists by Proposition 3.12 and Theorem 3.16. Using Proposition 3.13 we get that
+where $p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$ is as in Section 1. This integral exists by Propositions 3.12 and 3.16. Using Proposition 3.13 we get that
 $0$ $=$ $A_{r}+B_{r}$
@@
 for every $r\in\mathbb{N}$.
+The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as at the beginning, only now they are defined in an AB &\& SF way.
 We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded.
-The determination of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that we calculate semiformally. We define abbreviations
-$\mathrm{Exp}(-(x-i))$ $=$ $f((-1)x),\ f(x)=\mathrm{Exp}(x+(-i))\,\text{ and}$
-$\mathrm{Exp}(-((x-i)+i))$ $=$ $f((-1)x),\ f(x)=g(x+i),\ g(x)=\mathrm{Exp}(x+(-i))\,.$
-Then
+The computation of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that now we calculate semiformally.
 $B_{r}$ $=$ ${\textstyle\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{i}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)}$
@@
 […]
-Recall that $x$ is a formal variable. Using Proposition 3.12, Theorem 3.16, and the definition of the polynomial $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
-It remains to bound the numbers
+Using Propositions 3.12 and 3.16, and the definition of the polynomial $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
+It remains to bound, in an AB &\& SF way, the numbers
 $A_{r}=\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{0}^{i}p_{r}(x)\cdot\mathrm{Exp}(-x)\,.$}
@@
 $p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$
-has degree less than $(n+1)(r+1)$ and each coefficient of it is in absolute value at most $l^{r+1}$ where $l\in\mathbb{N}$ depends only on $n$.
+has degree less than $(n+1)(r+1)$ and each coefficient in it is in absolute value at most $l^{r+1}$ where $l\in\mathbb{N}$ depends only on $n$.
 By Propositions 3.8 and 3.15 each of the above $n+1$ Newton integrals in $A_{r}$ is in absolute value at most
 $(n+1)(r+1)\cdot l^{r+1}\cdot n^{(n+1)(r+1)}\cdot\mathrm{e}^{n}\,.$
-Thus $|A_{r}|\leq c^{r}$ for every $r\in\mathbb{N}$ and some constant $c\geq 1$. We get the same contradiction as in Section 1. $\Box$
-The previous proof is quasiformal and not formal because of Definitions 3.3 and 3.11. It clearly satisfies conditions (A) and (B).
+Thus $|A_{r}|\leq c^{r}$ for every $r\in\mathbb{N}$ and some constant $c\geq 1$. We get a contradiction as before. $\Box$
+The previous proof is semiformal and not formal due to the use of various semiformal operations and many values of formal power series. It is clearly an AB proof.

### v4

@@
 $:=$ ${\textstyle\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x+\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x}\,.$
-As for the quantity $A_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x$, it is easy to estimate it for $r\in\mathbb{N}$ by means of M-L bounds on integrals as
-$|A_{r}|\leq c^{r}$
+As for the quantity $A_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x$, it is easy to estimate it by means of M-L bounds on integrals as
+$|A_{r}|\leq c^{r},\ r\in\mathbb{N}\,$,
 where $c\geq 1$ is a constant (depending only on the coefficients $a_{0}$, $a_{1}$, $\dots$, $a_{n}$).
@@
 $f=f(x)=\sum_{n\geq 0}a_{n}x^{n}$}
-with $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also as $x^{n}]f$. The set $R[[x]]$ is endowed with the operations
+with $a_{n}=f(n)$.
+We denote the coefficient $a_{n}$ also as $x^{n}]f$.
+We may work with functionals on $R[[x]]$, which are maps from $R[[x]]$ to $R$.
+For example, for every $n\in\mathbb{N}_{0}$ we have the functional $f\mapsto[x^{n}]f$.
+Besides functionals, the set $R[[x]]$ is endowed with the operations of addition and multiplication
 ${\textstyle f(x)+g(x)=\sum_{n\geq 0}a_{n}x^{n}+\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}(a_{n}+b_{n})x^{n}\,\text{ and}}$
-${\textstyle f(x)\cdot g(x)=\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}\,,}$
-and neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. Then
+${\textstyle f(x)\cdot g(x)=\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}}$
+which have neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. For any $c,d\in R$ and $f,g\in R[[x]]$ we may form the linear combination
+$cf(x)+dg(x)=\sum_{n\geq 0}(ca_{n}+db_{n})x^{n}\,.$}
+This is in fact a family of binary operations on $R[[x]]$, one for every pair $c,d$. The structure
 R[[x]]=R[[x]], $\langle R[[x]],\,0,\,1$$,\cdot\rangle$
-is a domain and $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is the (formal) derivative,
+is a domain, in which $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is the (formal) derivative,
 $\big(\sum_{n\geq 0}a_{n}x^{n}\big)^{\prime}=\sum_{n\geq 0}(n+1)_{R}\cdot a_{n+1}x^{n}\,.$}
@@
 In other words, $f(x)\in K\{x\}$ iff $f(x)$ has an infinite radius of convergence.
-The binary and unary operations on $R[[x]]$ listed so far, ++, $\cdot$, $(\cdot)^{\prime}$, and $\int(\cdot)$, are formal in the sense that, for example, for every $n\in\mathbb{N}_{0}$ the coefficient
-$x^{n}](f\cdot g)$
-is determined by only finitely many coefficients in $f$ and $g$, and the same holds for the other operations.
-In formal proofs involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$, only formal operations may be used, and it is forbidden to use the functional values
+The functionals and operations on $R[[x]]$ we have listed so far, $f\mapsto[x^{n}]f$, ++, $\cdot$, $(\cdot)^{\prime}$, $\int(\cdot)$ and linear combinations, are formal in the sense that, for example, every coefficient
+$x^{n}](f\cdot g),\ n\in\mathbb{N}_{0}\,$,
+is determined by only finitely many coefficients in $f$ and $g$, and the same holds for the other functionals and operations.
+We say that a proof involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$ is formal if it employs only formal functionals and operations, and if it does not use any functional values
 $f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k},\ c\in K\,.$}
@@
 However, we can present two such proofs in the next two sections if formality is relaxed to semiformality.
-Semiformal operations (in Section 3 we introduce several of them) are defined on $K\{x\}$, may use limit transitions, and may use at most countably many values $f(c)$ for $f\in K[[x]]$ and $c\in K$.
-Semiformal proofs concerning elements in $K[[x]]$ are proofs that satisfy the following condition.
-Only formal and semiformal operations on $K[[x]]$ are used. For every formal power series involved in the proof, only at most countably many of its values are used.
+Semiformal functionals are defined on $K\{x\}$, may use limit transitions, and may use at most countably many values $f(c)$ for $f\in K[[x]]$ and $c\in K$.
+The same holds for semiformal operations.
+For examples of semiformal functionals and operations on $\mathbb{R}[[x]]$ see Definitions 3.7 and 3.3, respectively.
+We say that a proof involving elements in $K[[x]]$ is semiformal if it satisfies the following condition.
+Only formal and semiformal functionals and operations on $K[[x]]$ are used in the proof. For every formal power series involved in the proof, only at most countably many of its values are used.
 In Sections 2 and 3 we give two AB &\& SF proofs of the transcendence of $\mathrm{e}$.
@@
 We prove, with the help of Proposition 2.1 and Theorem 2.2, that the number $\mathrm{e}$ is transcendental.
-We assume, for the contradiction, that $b_{j},\alpha_{j}\in\mathbb{Z}$, $j\in[t]$ ($\{1,2,\dots,t\}$), are $2t$ integers such that every $b_{j}\neq 0$, the $\alpha_{j}$ are mutually distinct, and that
+We assume, for the contradiction, that $b_{j},\alpha_{j}\in\mathbb{Z}$, $j=1,2,\dots,t$, are $2t\geq 2$ integers such that every $b_{j}\neq 0$, the $\alpha_{j}$ are mutually distinct, and that
 $\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}}=0\,.$}
@@
 It is due to this author and first appeared in early versions of [7].
-The idea is to translate Hilbert’s proof, which we described at the beginning, into the language of formal power series.
+The idea is to translate Hilbert’s proof, described in Section 1, into the language of formal power series.
 We work in the domains of formal power series $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$.
-We introduce further operations on $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$ besides those we mentioned in Section 1.
+We introduce further functionals and operations on $\mathbb{R}[[x]]$ besides those mentioned in Section 1.
 For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
 $f(ax)=\sum_{n\geq 0}a^{n}a_{n}x^{n}\,.$}
-The operation $f(x)\mapsto f(ax)$, which we call the constant inner product, is formal. Clearly, for every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
+The family of operations
+$\{f(x)\mapsto f(ax)\colon\;a\in\mathbb{R}\}\,$,
+which we call constant inner products, consists of formal unary operations. For every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
 $(af+bg)(cx)=af(cx)+bg(cx)\,\text{ and }\,(fg)(cx)=f(cx)g(cx)\,.$
-If $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ then $f(0x)=a_{0}$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
+Also, $f(0x)=[x^{0}]f$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
 In the next proposition, whose straightforward proof we omit, we collect properties of the operations $f\mapsto f^{\prime}$ and $f\mapsto\int f$.
@@
 is called the (formal) exponential.
-It is easy to show that $\mathrm{Exp}\in\mathbb{R}\{x\}$, $\mathrm{Exp}^{\prime}=\mathrm{Exp}$ and that $\int\mathrm{Exp}=\mathrm{Exp}-1$.
-We introduce a semiformal operation.
-Definition 3.3 (semiformal shifts)
+Clearly, $\mathrm{Exp}\in\mathbb{R}\{x\}$, $\mathrm{Exp}^{\prime}=\mathrm{Exp}$ and $\int\mathrm{Exp}=\mathrm{Exp}-1$.
+We introduce a family of semiformal unary operations.
+Definition 3.3 (shifts)
 Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ be in $\mathbb{R}\{x\}$ and let $b\in\mathbb{R}$. We define the formal power series $f(x+b)$ in $\mathbb{R}[[x]]$ by
@@
 Clearly, $f(x+0)=f(x)$ and if $f(x)\in\mathbb{R}[x]$ then $f(x)\mapsto f(x+b)$ is the usual polynomial substitution.
-We combine the operations of constant inner product and semiformal shift by means of brackets.
+We combine the operations of constant inner product and shift by means of brackets.
 For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
-In the next two propositions, we show that semiformal shifts behave in desirable ways.
+We next show that shifts behave in desirable ways.
 Proposition 3.4
@@
 ${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{m=n}^{l}\binom{l-n}{l-m}c^{l-m}d^{m-n}=}$
-${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{j=0}^{l-n}\binom{l-n}{j}c^{j}d^{l-n-j}=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}=}$
-$=[x^{n}]f(x+(c+d))\,.$
+${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{j=0}^{l-n}\binom{l-n}{j}c^{j}d^{l-n-j}=}$
+${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}=[x^{n}]f(x+(c+d))\,.}$
 In the third line, we changed the order of summation. This is allowed because $f(x)$ is convergent. In the fourth line, we used the binomial theorem. $\Box$
@@
 $(f\cdot g)(x+c)=f(x+c)\cdot g(x+c)\;.$
-Proof.
-Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ and $c$ be as stated.
-The coefficient of $x^{n}$ in the formal power series $f(x+c)$, respectively $g(x+c)$, equals $\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}$, respectively $\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}$.
-Thus
+Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ and $c$ be as stated. Then the coefficients $x^{n}]f(x+c)$ and $x^{n}]g(x+c)$ equal to
+$\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}\,$,}
+respectively. Thus
 ${\textstyle[x^{n}]f(x+c)g(x+c)=\sum_{k=0}^{n}\big(\sum_{m\geq k}\binom{m}{k}a_{m}c^{m-k}\sum_{l\geq n-k}\binom{l}{n-k}b_{l}c^{l+k-n}\big)}$
@@
 ${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\sum_{k=0}^{n}\binom{m}{k}\binom{l}{n-k}a_{m}b_{l}c^{l+m-n}=}$
-${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\binom{m+l}{n}a_{m}b_{l}c^{l+m-n}=\sum_{j\geq n}\binom{j}{n}\big(\sum_{m=0}^{j}a_{m}b_{j-m}\big)c^{j-n}=}$
-$=[x^{n}](fg)(x+c)\,.$
+${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\binom{m+l}{n}a_{m}b_{l}c^{l+m-n}}$
+${\textstyle=\sum_{j\geq n}\binom{j}{n}\big(\sum_{m=0}^{j}a_{m}b_{j-m}\big)c^{j-n}=[x^{n}](fg)(x+c)\,.}$
 In the third line, we changed the order of summation. This is allowed since $f(x)$ is convergent. In the fourth line, we used Vandermonde’s identity ([9]). $\Box$
-The following semiformal exponential identity is somewhat counter-intuitive (at first we got it wrong).
+The following semiformal exponential identity is perhaps counter-intuitive; at first we got it wrong.
 Proposition 3.6
@@
 ${\textstyle f\big(x+\frac{b}{a}\big)}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}(a^{m}/m!)(b/a)^{m-n}\big)x^{n}}$
-$=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}=\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)(a^{n}/n!)x^{n}}$
+$=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}}$
+$=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)(a^{n}/n!)x^{n}}$
 $=$ $\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$
@@
 We will use this identity only with a=1a=-1.
-We introduce Newton integrals of convergent formal power series.
+We introduce a family of semiformal functionals on $\mathbb{R}[[x]]$.
 Definition 3.7 (Newton integral of fps)
@@
 By item 3 of Proposition 3.1, $\int f\in\mathbb{R}\{x\}$ and the integral is always defined.
-The Newton integral is a semiformal operation because it uses two values of a formal power series.
-We prove three integral transformations for it: linear combination, additive decomposition, and shift of the integration interval.
+We prove three transformations for it: linear combination, additive decomposition, and shift of the integration interval.
 The first two transformations are straightforward.
@@
 This is allowed since the involved series absolutely converge.
-In the fourth line, we add
+In the fourth line, we added
 $0=\sum_{m\geq 0}\frac{a_{m}}{m+1}\binom{m+1}{0}b^{m+1}v^{0}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\binom{m+1}{0}b^{m+1}u^{0}\,.$}
-In the fifth line, we use the binomial theorem. $\Box$
-We introduce improper Newton integrals of convergent formal power series.
+In the fifth line, we used the binomial theorem. $\Box$
+We introduce another family of semiformal functionals on $\mathbb{R}[[x]]$.
 Definition 3.11 (improper Newton $\int$ of fps)
@@
 We call $L$ the improper Newton integral of $f(x)$ from $u$ to $\infty$.
-We leave the proof of the independence of the limit $L$ on the sequences $(b_{n})$ to the reader as an exercise.
+We leave the proof of the independence of the limit $L$ on the sequence $(b_{n})$ to the reader as an exercise.
 We prove the same three integral transformations as before.
@@
 $\Box$
-We need to bound Newton integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)$ is in $\mathbb{Z}[x]$. We easily obtain the next bound based on two values of the primitive.
+We need bounds, obtained in an AB way, on integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)\in\mathbb{Z}[x]$.
+We obtain them via the next bound which is based on two values of the primitive.
 Proposition 3.15
@@
 $\Box$
-It is crucial that the bound is exponential and not factorial in $k$.
-The key step in Hilbert’s proof of the transcendence of $\mathrm{e}$ is Euler’s identity. We prove it in the language of improper Newton integrals of convergent formal power series.
+As we know, it is crucial that the bound is exponential and not factorial in $k$.
+The key step in Hilbert’s proof is Euler’s identity. We prove it now in an AB &\& SF way.
 Proposition 3.16
@@
 $\Box$
-We are ready to translate Hilbert’s proof of the transcendence of $\mathrm{e}$ into formal power series. We again assume for the contradiction that
+We are ready to state Hilbert’s proof in an AB &\& SF way. We again assume for the contradiction that
 $a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0$
@@
 for every $r\in\mathbb{N}$.
-The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as at the beginning, only now they are defined in an AB &\& SF way.
+The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as before, only now they are defined in an AB &\& SF way.
 We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded.
@@
 […]
-Using Propositions 3.12 and 3.16, and the definition of the polynomial $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
+Using Propositions 3.12 and 3.16, and the definition of $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
 It remains to bound, in an AB &\& SF way, the numbers
@@
 Thus $|A_{r}|\leq c^{r}$ for every $r\in\mathbb{N}$ and some constant $c\geq 1$. We get a contradiction as before. $\Box$
-The previous proof is semiformal and not formal due to the use of various semiformal operations and many values of formal power series. It is clearly an AB proof.
+The previous proof is semiformal and not formal due to the use of semiformal functionals and operations, and use of many values of formal power series. It is clearly an AB proof.

### v5

@@
 The second proof uses improper integrals of formal power series and is due to this author.
-We explain what “semiformal” means and indicate yet another analytical proof of the transcendence of $\mathrm{e}$.
-Like the two algebraic proofs, this proof avoids uncountable sets.
-It is due to this author in [7].
+We explain what “semiformal” means.
 1 Hilbert’s proof that $\mathrm{e}$ is transcendental. Why this proof is not completely satisfactory. How to fix it
@@
 Hilbert’s proof is elegant and short, but it has the unsatisfactory feature that it employs uncountable sets, for example, the integrands in Euler’s identity.
-We explain how to improve upon it by proving the transcendence of $\mathrm{e}$ so that two conditions hold.
+We explain how to improve upon it by proving the transcendence of $\mathrm{e}$ in a way satisfying two conditions.
 Only hereditarily at most countable (HMC) sets are directly used in the proof.
@@
 the set $x_{n}$ is at most countable (finite or countable).
-As for uses of mathematical logic in analysis, we quote a referee of [7]: “That basic real analysis is essentially a countable endeavor has been well-known for over 100 years.
-My secondhand understanding is that Weyl articulated this very well in Das Kontinuum, which nowadays we would say developed analysis in $\mathrm{ACA}_{0}$”.
-The reader can check [8] for $\mathrm{ACA}_{0}$.
-Our more ambitious and demanding goal is to obtain an AB proof of the transcendence of $\mathrm{e}$.
 In a related result, Kaibkhanov and Skopenkov [6] gave an AB proof of the transcendence of the Mahler number $\sum_{n\geq 0}2^{-2^{n}}$ based on binary expansions.
-There are at least three ways to such proof.
-The first way is to replace real functions of the type $f\colon[a,b]\to\mathbb{R}$, where $a<b$ are real numbers and $\mathbb{R}$ is the set of real numbers, with the functions
-$f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}\,$,
-where $\mathbb{Q}$ is the set of fractions (if real numbers are Dedekind cuts, every such function $f$ is an HMC set), and to develop a sufficient fragment of real analysis for such functions.
-We realized this way in [7].
-Is there any difference in the treatment of the two kinds of functions?
-The above mentioned referee thinks not: “Continuous functions on the reals are determined by their values on the rationals, so everything works out in the usual way.
-I don’t see how what the article is doing is so very different from developing analysis in (full) second-order arithmetic”.
-An obvious meta-argument refutes the first claim: if it were the case that “everything works out in the usual way”, then some enthusiasts would be teaching courses in analysis based on functions of the type $f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}$ for years.
-Since you have not heard about such courses, there must be some difficulties with this approach.
-Indeed, there are, but they can be overcome.
-The interested reader is referred to [7] for details.
-The second claim is refuted by the fact that in [7] we provide an AB proof.
-We do not use any technical notion of mathematical logic, and our approach in [7] is therefore very different from the countable treatments of real analysis offered by mathematical logic.
-For examples of the latter approaches, see Weaver [10, 11].
-This analytical digression places the algebraic results of the present article in context.
-These results are two AB proofs of the transcendence of $\mathrm{e}$.
-Both are based on formal power series.
+There are at least four ways to an AB proof of the transcendence of $\mathrm{e}$.
+For the first two see [7].
+The other two ways are the topic of the present article.
+They are two AB proofs of the transcendence of $\mathrm{e}$ based on formal power series.
 The first proof in Section 2 is a specialization of the proof in [2].
@@
 It is clear that the set $\mathbb{R}$ is mentioned only for convenience and can be easily formulated away.
-We allow such uses of non-HMC sets in AB proofs for the sake of natural and easily legible formulations.
+We allow such mentions of non-HMC sets in AB proofs for the sake of natural and legible formulations.
 In conclusion of the introduction, we review formal power series and explain the meaning of the adjective “semiformal”.
@@
 We denote the coefficient $a_{n}$ also as $x^{n}]f$.
-We may work with functionals on $R[[x]]$, which are maps from $R[[x]]$ to $R$.
+We work with functionals on $R[[x]]$, which are maps from $R[[x]]$ to $R$.
 For example, for every $n\in\mathbb{N}_{0}$ we have the functional $f\mapsto[x^{n}]f$.
@@
 ${\textstyle f(x)\cdot g(x)=\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}}$
-which have neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. For any $c,d\in R$ and $f,g\in R[[x]]$ we may form the linear combination
+which have neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. For any $c,d\in R$ and $f,g\in R[[x]]$ we form the linear combination
 $cf(x)+dg(x)=\sum_{n\geq 0}(ca_{n}+db_{n})x^{n}\,.$}
@@
 $\lim_{n\to\infty}a_{n}c^{n}=0_{K}\,.$
-In other words, $f(x)\in K\{x\}$ iff $f(x)$ has an infinite radius of convergence.
-The functionals and operations on $R[[x]]$ we have listed so far, $f\mapsto[x^{n}]f$, ++, $\cdot$, $(\cdot)^{\prime}$, $\int(\cdot)$ and linear combinations, are formal in the sense that, for example, every coefficient
-$x^{n}](f\cdot g),\ n\in\mathbb{N}_{0}\,$,
+In other words, $f(x)\in K\{x\}$ iff $f(x)$ has infinite radius of convergence.
+The functionals and operations on $R[[x]]$ we have listed so far, $f\mapsto[x^{n}]f$, ++, $\cdot$, $(\cdot)^{\prime}$, $\int(\cdot)$ and linear combinations, are formal in the sense that, for example, every coefficient in the product $f\cdot g$,
+$[x^{n}](f\cdot g)=\sum_{j=0}^{n}[x^{j}]f\cdot[x^{n-j}]g\,$,}
 is determined by only finitely many coefficients in $f$ and $g$, and the same holds for the other functionals and operations.
-We say that a proof involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$ is formal if it employs only formal functionals and operations, and if it does not use any functional values
-$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k},\ c\in K\,.$}
+We say that a proof involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$ is formal if it employs only formal functionals and operations, and if it does not use any values of $f(x)$ ($c\in K$)
+$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k}\,.$}
 It is an open problem to give a completely formal AB proof of the transcendence of $\mathrm{e}$ based on formal power series. No such proof is known to this author.
@@
 and $k!\,|\,v_{n}(k)$ because every coefficient in $W(x)$ is divisible by $k!$.
-We prove the claim: for any $r$ with $0\leq r\leq k-1$ we indeed have, as $\binom{n}{r}=(-1)^{n-r}\binom{-r-1}{n-r}$ for $n\geq r$, that
+We prove the claim: for every $r$ with $0\leq r\leq k-1$ we indeed have, since $\binom{n}{r}=(-1)^{n-r}\binom{-r-1}{n-r}$ for $n\geq r$, that
 ${\textstyle\sum_{n\geq 0}(n)_{r}u_{n-r}x^{n}}$ $=$ ${\textstyle\sum_{j=1}^{t}r!\cdot b_{j}x^{r}\sum_{n\geq r}\binom{n}{r}\alpha_{j}^{n-r}x^{n-r}}$
@@
 Although the article [2] of Beukers, Bézivin, and Robba is a pioneering one, as it is the first AB proof of the transcendence of the number $\mathrm{e}$ (in fact, of course, of the full LW theorem), it is little known; in the database of Mathematical Reviews, it currently has only 11 citation (January 2026).
-3 The proof by formal integration
+3 The proof by semiformal integration
 We present the second AB &\& SF proof of the transcendence of $\mathrm{e}$.
@@
 For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
-We next show that shifts behave in desirable ways.
+We show that shifts behave in desirable ways.
 Proposition 3.4
@@
 $=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}}$
-$=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)(a^{n}/n!)x^{n}}$
-$=$ $\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$
+$=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)(a^{n}/n!)x^{n}}=\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$
 $\Box$
-We will use this identity only with a=1a=-1.
+We use this identity only for a=1a=-1.
 We introduce a family of semiformal functionals on $\mathbb{R}[[x]]$.
@@
 By item 3 of Proposition 3.1, $\int f\in\mathbb{R}\{x\}$ and the integral is always defined.
-We prove three transformations for it: linear combination, additive decomposition, and shift of the integration interval.
+We obtain three transformations for it: linear combination, additive decomposition, and shift of the integration interval.
 The first two transformations are straightforward.
@@
 $\int_{u}^{v}(af+bg)=a\int_{u}^{v}f+b\int_{u}^{v}g\,.$}
-Proof. This follows from the equality $\int(af+bg)=a\int f+b\int g$ in Proposition 3.1. $\Box$
+Proof. This follows from the equality $\int(af+bg)=a\int f+b\int g$ in item 2 of Proposition 3.1. $\Box$
 Proposition 3.9
@@
 We leave the proof of the independence of the limit $L$ on the sequence $(b_{n})$ to the reader as an exercise.
-We prove the same three integral transformations as before.
+We obtain for the improper integral the same three integral transformations as before.
 Proposition 3.12
@@
 Let $f$, $u$ and $v$ be as stated, let $(b_{n})$ be any real sequence going to $\infty$ and let us suppose that the former improper integral exists.
-Thus the limit $L=\lim_{n\to\infty}\big(\int f\big)(b_{n})$ exists.
+Thus the limit $L=\lim_{n\to\infty}\big(\int f\big)(b_{n})$ ($\in\mathbb{R}$) exists.
 But then
@@
 ${\textstyle=\int_{u}^{v}f+\int_{v}^{+\infty}f\,.}$
-With the assumption of existence of $\int_{v}^{+\infty}f$ we argue similarly. $\Box$
+In the other way we argue similarly. $\Box$
 The third transformation is immediate as well.
@@
 The key step in Hilbert’s proof is Euler’s identity. We prove it now in an AB &\& SF way.
-Proposition 3.16
+Proposition 3.16 (Euler’s identity)
 For every $k\in\mathbb{N}_{0}$, the integral
@@
 and $F_{k}=k\cdot F_{k-1}-F_{k}^{\prime}$. By item 4 of Proposition 3.1, $\int F_{k}^{\prime}=F_{k}-[x^{0}]F_{k}=F_{k}$. By Proposition 3.12 and induction,
-$I_{k}$ $=$ $\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k}}=k\cdot\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k-1}-}\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k}^{\prime}}$
+$I_{k}$ $=$ $\lim_{n\to\infty}{\textstyle\int_{0}^{a_{n}}F_{k}}=k\cdot\lim_{n\to\infty}{\textstyle\int_{0}^{a_{n}}F_{k-1}-}\lim_{n\to\infty}{\textstyle\int_{0}^{a_{n}}F_{k}^{\prime}}$
 $=$ $k\cdot I_{k-1}-\lim_{n\to\infty}\big(F_{k}(a_{n})-F_{k}(0)\big)$
@@
 […]
-Using Propositions 3.12 and 3.16, and the definition of $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
+Using Propositions 3.12 and 3.16, and the definition of $p_{r}(x)$, we obtain the above properties of $B_{r}$.
 It remains to bound, in an AB &\& SF way, the numbers
@@
 Thus $|A_{r}|\leq c^{r}$ for every $r\in\mathbb{N}$ and some constant $c\geq 1$. We get a contradiction as before. $\Box$
-The previous proof is semiformal and not formal due to the use of semiformal functionals and operations, and use of many values of formal power series. It is clearly an AB proof.
+The previous proof is semiformal and not formal due to the use of semiformal functionals and operations, and the use of many values of formal power series. Clearly it is an AB proof.

### v6

@@
 We remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert.
-Then, using formal power series, we give two algebraic semiformal proofs of this result.
-The first proof is a specialization of the proof of the Lindemann–Weierstrass theorem found by Beukers, Bézivin and Robba [2].
+Using formal power series, we then give two algebraic proofs of this result.
+The first proof is a specialization of the proof of Beukers, Bézivin and Robba of the Lindemann–Weierstrass theorem.
 The second proof uses improper integrals of formal power series and is due to this author.
-We explain what “semiformal” means.
-1 Hilbert’s proof that $\mathrm{e}$ is transcendental. Why this proof is not completely satisfactory. How to fix it
+We explain in what respect both proofs improve upon Hilbert’s proof.
+1 Hilbert’s proof and how to improve it
 The transcendence of the number $\mathrm{e}$ was first proven by Hermite [4].
 Hilbert’s proof [5] goes as follows.
-We suppose for the contradiction that there exists $n+1$, $n\in\mathbb{N}$ ($\{1,2,\dots\}$), integers $a_{0}$, $a_{1}$, $\dots$, $a_{n}$, not all of them 0, such that
+We suppose for the contradiction that there exists $n+1$ integers $a_{0}$, $a_{1}$, $\dots$, $a_{n}$, not all of them 0, such that
 $P:=a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0\,.$
-We may assume that $a_{0}\neq 0$. For $r\in\mathbb{N}$ we define integral polynomials
+We may assume that $a_{0}\neq 0$. For any $r\in\mathbb{N}$ ($\{1,2,\dots\}$) we define integral polynomials
 $p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}\,$,
@@
 $\Box$
-Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof using only proper integrals $\int_{a}^{b}$. A similar proof is given in Conrad [3].
-Hilbert’s proof is elegant and short, but it has the unsatisfactory feature that it employs uncountable sets, for example, the integrands in Euler’s identity.
-We explain how to improve upon it by proving the transcendence of $\mathrm{e}$ in a way satisfying two conditions.
-Only hereditarily at most countable (HMC) sets are directly used in the proof.
-No technical notion of mathematical logic such as, for example, the notion of a formula, is used in the proof.
-We call proofs satisfying these two conditions AB proofs. A set $x$ is HMC if for every index $n\in\mathbb{N}_{0}$ and every chain of sets
-$x_{n}\in x_{n-1}\in\dots\in x_{0}=x\,$,
-the set $x_{n}$ is at most countable (finite or countable).
-In a related result, Kaibkhanov and Skopenkov [6] gave an AB proof of the transcendence of the Mahler number $\sum_{n\geq 0}2^{-2^{n}}$ based on binary expansions.
-There are at least four ways to an AB proof of the transcendence of $\mathrm{e}$.
-For the first two see [7].
-The other two ways are the topic of the present article.
-They are two AB proofs of the transcendence of $\mathrm{e}$ based on formal power series.
-The first proof in Section 2 is a specialization of the proof in [2].
-The second proof in Section 3 is due to this author.
-We presented it already in early versions of [7], and we rework it here.
-The word “directly” in condition (A) means that non-HMC sets may appear in the proof only as supersets of the main, directly used sets, which have to be HMC.
-We may say, for example, that for every two sequences $(a_{n}),(b_{n})\subset\mathbb{R}$ we have $\dots\;$.
-It is clear that the set $\mathbb{R}$ is mentioned only for convenience and can be easily formulated away.
-We allow such mentions of non-HMC sets in AB proofs for the sake of natural and legible formulations.
-In conclusion of the introduction, we review formal power series and explain the meaning of the adjective “semiformal”.
-If $R$ is an (integral) domain, we write $R[[x]]$ for the uncountable set of maps $f\colon\mathbb{N}_{0}\to R$.
-They are written as infinite formal sums
-$f=f(x)=\sum_{n\geq 0}a_{n}x^{n}$}
-with $a_{n}=f(n)$.
-We denote the coefficient $a_{n}$ also as $x^{n}]f$.
-We work with functionals on $R[[x]]$, which are maps from $R[[x]]$ to $R$.
-For example, for every $n\in\mathbb{N}_{0}$ we have the functional $f\mapsto[x^{n}]f$.
-Besides functionals, the set $R[[x]]$ is endowed with the operations of addition and multiplication
-${\textstyle f(x)+g(x)=\sum_{n\geq 0}a_{n}x^{n}+\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}(a_{n}+b_{n})x^{n}\,\text{ and}}$
-${\textstyle f(x)\cdot g(x)=\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}}$
-which have neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. For any $c,d\in R$ and $f,g\in R[[x]]$ we form the linear combination
-$cf(x)+dg(x)=\sum_{n\geq 0}(ca_{n}+db_{n})x^{n}\,.$}
-This is in fact a family of binary operations on $R[[x]]$, one for every pair $c,d$. The structure
-R[[x]]=R[[x]], $\langle R[[x]],\,0,\,1$$,\cdot\rangle$
-is a domain, in which $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is the (formal) derivative,
-$\big(\sum_{n\geq 0}a_{n}x^{n}\big)^{\prime}=\sum_{n\geq 0}(n+1)_{R}\cdot a_{n+1}x^{n}\,.$}
-Less often one uses the (formal) primitive
-$\int\sum_{n\geq 0}a_{n}x^{n}=\sum_{n\geq 0}\frac{a_{n}}{n+1}x^{n+1}=\sum_{n\geq 1}\frac{a_{n-1}}{n}x^{n}\,.$}
-In this article, $\int$ is more important than the derivative.
-If $K$ is the normed field $K=\mathbb{R}$ or $K=\mathbb{C}$ (complex numbers), we denote by $K\{x\}$ the subdomain of $K[[x]]$ formed by the convergent formal power series $f(x)=\sum_{n\geq 0}a_{n}x_{n}$.
-They satisfy for every $c\in K$ that
-$\lim_{n\to\infty}a_{n}c^{n}=0_{K}\,.$
-In other words, $f(x)\in K\{x\}$ iff $f(x)$ has infinite radius of convergence.
-The functionals and operations on $R[[x]]$ we have listed so far, $f\mapsto[x^{n}]f$, ++, $\cdot$, $(\cdot)^{\prime}$, $\int(\cdot)$ and linear combinations, are formal in the sense that, for example, every coefficient in the product $f\cdot g$,
-$[x^{n}](f\cdot g)=\sum_{j=0}^{n}[x^{j}]f\cdot[x^{n-j}]g\,$,}
-is determined by only finitely many coefficients in $f$ and $g$, and the same holds for the other functionals and operations.
-We say that a proof involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$ is formal if it employs only formal functionals and operations, and if it does not use any values of $f(x)$ ($c\in K$)
-$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k}\,.$}
-It is an open problem to give a completely formal AB proof of the transcendence of $\mathrm{e}$ based on formal power series. No such proof is known to this author.
-However, we can present two such proofs in the next two sections if formality is relaxed to semiformality.
-Semiformal functionals are defined on $K\{x\}$, may use limit transitions, and may use at most countably many values $f(c)$ for $f\in K[[x]]$ and $c\in K$.
-The same holds for semiformal operations.
-For examples of semiformal functionals and operations on $\mathbb{R}[[x]]$ see Definitions 3.7 and 3.3, respectively.
-We say that a proof involving elements in $K[[x]]$ is semiformal if it satisfies the following condition.
-Only formal and semiformal functionals and operations on $K[[x]]$ are used in the proof. For every formal power series involved in the proof, only at most countably many of its values are used.
-In Sections 2 and 3 we give two AB &\& SF proofs of the transcendence of $\mathrm{e}$.
+Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof using only proper integrals $\int_{a}^{b}$. A similar proof appears in Conrad [3].
+Hilbert’s proof is a mathematical gem, but there is still a large room for improvement.
+The functions $x^{n}\mathrm{e}^{-x}\colon[0,+\infty)\to\mathbb{R}$, which play a key role in the proof, are uncountable sets.
+Is it possible to give a proof of the transcendence of $\mathrm{e}$ without substantially using uncountable sets?
+(For example, when one proves the formula $\sum_{n\geq 0}a^{n}=\frac{1}{1-a}$ for $a\in(-1,1)$, one produces such a proof, even though the uncountable set (1,1)(-1,1) is mentioned; it is clear how to formulate (1,1)(-1,1) away.
+Of course, individual real numbers have to be represented by Dedekind’s cuts and not by Cantor’s equivalence classes of Cauchy sequences of fractions.) Yes, this is possible, such a proof was found already in 1990 by Beukers, Bézivin and Robba [2]; they actually proved in this way a much more general result, the Lindemann–Weierstrass theorem.
+What are the main countable players in their proof which replace the functions $x^{n}\mathrm{e}^{-x}$?
+The complex sequences
+$(a_{n})=(a_{0},\,a_{1},\,a_{2},\,\dots),\ a_{n}\in\mathbb{C}\,$,
+known also as formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with complex coefficients. In contrast to functions $x^{n}\mathrm{e}^{-x}$, the variable $x$ is now only formal.
+Why do we hear almost nothing about [2]?
+(In March 2026 it had $1$ citation in Mathematical Reviews.) In the approach of eliminating uncountable sets from proofs, [2] is a remarkable achievement, even if the authors are probably not aware of it (that approach is not mentioned in [2]).
+However, only a small minority of professional mathematicians, to which this author belongs, care about the distinction of using or not using uncountable sets in proofs; see, for example, the forthcoming book [7].
+The proof of Beukers, Bézivin and Robba deserves a new edition, so to say, and we review it in Section 2, but only in the specialization to the proof of the transcendence of $\mathrm{e}$.
+In Section 3 we present another uncountable-sets-free proof of the transcendence of $\mathrm{e}$ using formal power series.
+It was found by this author and is based on the natural idea of translating Hilbert’s proof to the language of formal power series.
+To conclude the Introduction, we reiterate that the main contribution of the two proofs in the next two sections is not that they use formal power series; the main novelty is that they avoid (substantial use of) uncountable sets.
 2 Specializing the proof of Beukers, Bézivin and Robba
@@
 $b_{1}\mathrm{e}^{\alpha_{1}}+b_{2}\mathrm{e}^{\alpha_{2}}+\dots+b_{t}\mathrm{e}^{\alpha_{t}}\neq 0$
-whenever $b_{j}$ and $\alpha_{j}$ are $2t$, with $t\in\mathbb{N}$, algebraic numbers in $\mathbb{C}$ such that every $b_{j}\neq 0$ and the $\alpha_{j}$ are mutually distinct.
+whenever $b_{j}$ and $\alpha_{j}$ are $2t\geq 2$ algebraic numbers in $\mathbb{C}$ such that every $b_{j}\neq 0$ and the $\alpha_{j}$ are mutually distinct.
 See Baker [1, Theorem 1.4 on pp.
@@
 In this section, we specialize their proof to $b_{j},\alpha_{j}\in\mathbb{Z}$, which gives the transcendence of $\mathrm{e}$.
-In order to clearly present the logical structure of their proof, we reorganize the exposition in [2].
+In order to clearly present the logical structure of the proof, we reorganize the exposition in [2].
 We first recall that a formal power series $f(x)$ in $\mathbb{C}[[x]]$ is rational if there exist two polynomials $p(x),q(x)\in\mathbb{C}[x]$ such that $q(0)=1$ and $q(x)f(x)=p(x)$; we may assume that $p(x)$ and $q(x)$ are coprime.
@@
 contradicts Proposition 2.1 because by Theorem 2.2 the formal power series $V(x)$ is rational. $\Box$
-We repeat that the proof in [2] is more general as it establishes the full LW theorem.
-In that proof, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of.
-The previous proof works with formal power series, but it is semiformal and not formal because the key player is the infinite sum
-$\sum_{n\geq 0}\frac{u_{n}}{n!}=0$}
-which is the value $f(1)$ of the formal power series $f(x)=\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}$. In fact, this is the only non-formal component in an otherwise formal proof.
-It is clear that the previous proof satisfies condition (A), which states that all directly used sets are HMC.
-It works directly only with complex formal power series, which are maps from $\mathbb{N}_{0}$ to $\mathbb{C}$.
-As explained in Section 1, non-HMC sets like $\mathbb{C}[[x]]$ are not really needed in the proof.
-Condition (B), which states that no formalism of mathematical logic is used, is satisfied as well.
-Although the article [2] of Beukers, Bézivin, and Robba is a pioneering one, as it is the first AB proof of the transcendence of the number $\mathrm{e}$ (in fact, of course, of the full LW theorem), it is little known; in the database of Mathematical Reviews, it currently has only 11 citation (January 2026).
+In the original proof, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of.
 3 The proof by semiformal integration
-We present the second AB &\& SF proof of the transcendence of $\mathrm{e}$.
-It is due to this author and first appeared in early versions of [7].
-The idea is to translate Hilbert’s proof, described in Section 1, into the language of formal power series.
-We work in the domains of formal power series $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$.
-We introduce further functionals and operations on $\mathbb{R}[[x]]$ besides those mentioned in Section 1.
+We work in the domains of formal power series $\mathbb{R}[[x]]$ and convergent formal power series $\mathbb{R}\{x\}$; the latter are formal power series with an infinite radius of convergence.
+For $f(x)\in\mathbb{R}[[x]]$ and $n\in\mathbb{N}_{0}$ we denote by $x^{n}]f(x)$ the coefficient of $x^{n}$ in $f(x)$.
 For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
@@
 Also, $f(0x)=[x^{0}]f$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
-In the next proposition, whose straightforward proof we omit, we collect properties of the operations $f\mapsto f^{\prime}$ and $f\mapsto\int f$.
+In the next proposition, whose straightforward proof we omit, we collect properties of the unary operations $f\mapsto f^{\prime}$ and $f\mapsto\int f$ on $\mathbb{R}[[x]]$. Recall that
+$\int\sum_{n\geq 0}a_{n}x^{n}=\sum_{n\geq 0}\frac{a_{n}}{n+1}x^{n+1}=\sum_{n\geq 1}\frac{a_{n-1}}{n}x^{n}\,.$}
 Proposition 3.1
@@
 $f(x+b)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)x^{n}\;.$}
+We call these operations on $\mathbb{R}[[x]]$ semiformal because coefficients in them arise by limit transitions; in the standard formal operations they are obtained by finite expressions.
 The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent.
@@
 $\Box$
-We need bounds, obtained in an AB way, on integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)\in\mathbb{Z}[x]$.
-We obtain them via the next bound which is based on two values of the primitive.
+We need bounds on integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)\in\mathbb{Z}[x]$. We obtain them via the next bound which is based on two values of the primitive.
 Proposition 3.15
@@
 As we know, it is crucial that the bound is exponential and not factorial in $k$.
-The key step in Hilbert’s proof is Euler’s identity. We prove it now in an AB &\& SF way.
+We prove Euler’s identity without using uncountable sets.
 Proposition 3.16 (Euler’s identity)
@@
 $\Box$
-We are ready to state Hilbert’s proof in an AB &\& SF way. We again assume for the contradiction that
+We are ready to state Hilbert’s proof in an uncountable-sets-free way. We again assume for the contradiction that
 $a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0$
@@
 for every $r\in\mathbb{N}$.
-The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as before, only now they are defined in an AB &\& SF way.
+The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as before, only now they are defined in an uncountable-sets-free way.
 We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded.
-The computation of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that now we calculate semiformally.
+The computation of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that now we compute semiformally.
 $B_{r}$ $=$ ${\textstyle\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{i}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)}$
@@
 Using Propositions 3.12 and 3.16, and the definition of $p_{r}(x)$, we obtain the above properties of $B_{r}$.
-It remains to bound, in an AB &\& SF way, the numbers
+It remains to bound, in an uncountable-sets-free way, the numbers
 $A_{r}=\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{0}^{i}p_{r}(x)\cdot\mathrm{Exp}(-x)\,.$}
@@
 Thus $|A_{r}|\leq c^{r}$ for every $r\in\mathbb{N}$ and some constant $c\geq 1$. We get a contradiction as before. $\Box$
-The previous proof is semiformal and not formal due to the use of semiformal functionals and operations, and the use of many values of formal power series. Clearly it is an AB proof.

### v7

@@
 In the third line, we changed the variable to y=xjy=x-j.
-In the fourth line, we used the above generalization of Euler’s identity and the expansions of $p_{r}(x+j)$, $j=0,1,\dots,n$.
+In the fourth line, we used the above generalization of Euler’s identity and the expansions of $p_{r}(x+j)$ for $j=0,1,\dots,n$.
 It is clear that $B_{r}\in\mathbb{Z}$ and is divisible by $r!$.
@@
 $\Box$
-Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof using only proper integrals $\int_{a}^{b}$. A similar proof appears in Conrad [3].
+Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof using only proper integrals $\int_{a}^{b}$. Similar proof appears in Conrad [3].
 Hilbert’s proof is a mathematical gem, but there is still a large room for improvement.
-The functions $x^{n}\mathrm{e}^{-x}\colon[0,+\infty)\to\mathbb{R}$, which play a key role in the proof, are uncountable sets.
-Is it possible to give a proof of the transcendence of $\mathrm{e}$ without substantially using uncountable sets?
+The functions $x^{n}\mathrm{e}^{-x}\colon[0,+\infty)\to\mathbb{R}$ in Euler’s identity, which are crucial in the proof, are uncountable sets.
+Is it possible to give a proof of the transcendence of $\mathrm{e}$ that avoids substantial use of uncountable sets?
 (For example, when one proves the formula $\sum_{n\geq 0}a^{n}=\frac{1}{1-a}$ for $a\in(-1,1)$, one produces such a proof, even though the uncountable set (1,1)(-1,1) is mentioned; it is clear how to formulate (1,1)(-1,1) away.
-Of course, individual real numbers have to be represented by Dedekind’s cuts and not by Cantor’s equivalence classes of Cauchy sequences of fractions.) Yes, this is possible, such a proof was found already in 1990 by Beukers, Bézivin and Robba [2]; they actually proved in this way a much more general result, the Lindemann–Weierstrass theorem.
-What are the main countable players in their proof which replace the functions $x^{n}\mathrm{e}^{-x}$?
+Of course, individual real numbers have to be represented by Dedekind’s cuts and not by Cantor’s equivalence classes of Cauchy sequences of fractions.) Yes, this is possible; such a proof was found already in 1990 by Beukers, Bézivin and Robba [2].
+They proved in this way a much more general result, the Lindemann–Weierstrass theorem.
+What are the main countable players in their proof replacing the functions $x^{n}\mathrm{e}^{-x}$?
 The complex sequences
 $(a_{n})=(a_{0},\,a_{1},\,a_{2},\,\dots),\ a_{n}\in\mathbb{C}\,$,
-known also as formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with complex coefficients. In contrast to functions $x^{n}\mathrm{e}^{-x}$, the variable $x$ is now only formal.
-Why do we hear almost nothing about [2]?
-(In March 2026 it had $1$ citation in Mathematical Reviews.) In the approach of eliminating uncountable sets from proofs, [2] is a remarkable achievement, even if the authors are probably not aware of it (that approach is not mentioned in [2]).
-However, only a small minority of professional mathematicians, to which this author belongs, care about the distinction of using or not using uncountable sets in proofs; see, for example, the forthcoming book [7].
-The proof of Beukers, Bézivin and Robba deserves a new edition, so to say, and we review it in Section 2, but only in the specialization to the proof of the transcendence of $\mathrm{e}$.
-In Section 3 we present another uncountable-sets-free proof of the transcendence of $\mathrm{e}$ using formal power series.
-It was found by this author and is based on the natural idea of translating Hilbert’s proof to the language of formal power series.
-To conclude the Introduction, we reiterate that the main contribution of the two proofs in the next two sections is not that they use formal power series; the main novelty is that they avoid (substantial use of) uncountable sets.
-2 Specializing the proof of Beukers, Bézivin and Robba
+known also as formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with complex coefficients.
+In contrast to functions $x^{n}\mathrm{e}^{-x}$, the variable $x$ is now only formal.
+A nice survey of applications of formal power series is Sambale [6].
+The article [2] is little known; in March 2026, it had $1$ citation in Mathematical Reviews.
+In the perspective of eliminating uncountable sets from proofs, [2] is a remarkable achievement.
+However, only a small minority of mathematicians, to which this author belongs, care about using or not using uncountable sets in proofs; see, for example, the forthcoming book [8].
+The proof of Beukers, Bézivin and Robba deserves a new edition, so to say.
+We review it in Section 2, but only in the specialization to the proof of the transcendence of $\mathrm{e}$.
+In Section 3 we present our uncountable-sets-free proof of the transcendence of $\mathrm{e}$.
+It is based on the translation of Hilbert’s proof to the language of formal power series.
+In conclusion, we reiterate that the main contribution of the two proofs in the next two sections is not that they use formal power series; the main novelty is that they avoid (substantial use of) uncountable sets.
+2 Specializing the proof of Beukers, Bézivin and Robba in [2]
 The Lindemann–Weierstrass theorem says that
@@
 where $r(x)\in\mathbb{C}[x]$, $t\in\mathbb{N}_{0}$, $m_{j}\in\mathbb{N}$, $\beta_{j,i},\alpha_{j}\in\mathbb{C}$ with $\beta_{j,m_{j}}\neq 0$, and the $\alpha_{j}$ are nonzero and mutually distinct.
-For $t=0$, the double sum is interpreted as 0, and the numbers $m_{j}$ and $\alpha_{j}$ come from the factorization
+For $t=0$, the double sum is defined as 0.
+The numbers $t$, $m_{j}$, and $\alpha_{j}$ come from the factorization
 $q(x)=\prod_{j=1}^{t}(1-\alpha_{j}x)^{m_{j}}\,.$}
-The partial fractions are the formal power series
+The partial fractions are then the formal power series
 $\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}=\sum_{n\geq 0}\binom{-i}{n}\beta_{j,i}\alpha_{j}^{n}\cdot x^{n}\,.$}
@@
 $v_{n}(k)=[x^{n}]q(x)V(x)=[x^{n}]q(x)W(x)$
-and $k!\,|\,v_{n}(k)$ because every coefficient in $W(x)$ is divisible by $k!$.
+(here $x^{n}]f(x)$ denotes the coefficient of $x^{n}$ in $f(x)$) and $k!\,|\,v_{n}(k)$ because every coefficient in $W(x)$ is divisible by $k!$.
 We prove the claim: for every $r$ with $0\leq r\leq k-1$ we indeed have, since $\binom{n}{r}=(-1)^{n-r}\binom{-r-1}{n-r}$ for $n\geq r$, that
@@
 Thus $\langle k,n\rangle\in N$ $\Rightarrow$ $v_{n}(k)=0$.
-For $\langle k,n\rangle\in\mathbb{N}^{2}$ we consider the norm $\|\langle k,n\rangle\|=n-2tk$ ($\in\mathbb{Z}$).
-We name the ordered pairs in the modified recurrence:
+For pairs $\langle k,n\rangle\in\mathbb{N}^{2}$ we consider the norm $\|\langle k,n\rangle\|=n-2tk$ ($\in\mathbb{Z}$).
+We name the pairs in the modified recurrence:
 $\gamma=\langle k,\,n\rangle,\,\beta_{0}=\langle k+1,\,n\rangle,\,\beta_{1}=\langle k,\,n-1\rangle,\,\dots,\,\beta_{t}=\langle k,\,n-t\rangle\,.$
@@
     For every $i=0,1,\dots,t$, the norm $\|\beta_{i}\|<\|\gamma\|$.
-From these three properties, from the property of the set $N$ and from the modified recurrence, it follows by induction on the norm that $v_{n}(k)=0$ for every $\langle k,n\rangle\in M$.
+From these three properties, from the implication for the set $N$ and from the modified recurrence, it follows by induction on the norm that $v_{n}(k)=0$ for every $\langle k,n\rangle\in M$.
 This concludes the proof because we see that every number $k\geq k_{0}$ has the required property.
@@
 We may assume that $\alpha_{j}\in\mathbb{N}$. We set $A:=\max(\{\alpha_{1},\dots,\alpha_{t}\})$ and define $u_{n}$ and $v_{n}$ as in Theorem 2.2. Then
-$\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\sum_{n\geq 0}\frac{\alpha_{j}^{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}x}$}
+$\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\sum_{n\geq 0}\frac{\alpha_{j}^{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}x}\ \ (\in\mathbb{C}[[x]])$}
 and
@@
 Let $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ ($\in\mathbb{Z}[[x]]$).
-As we know from the previous proof, $v_{n}-nv_{n-1}=u_{n}$ for $n\in\mathbb{N}$ and $v_{0}=u_{0}$.
+We know from the proof of Theorem 2.2 that $v_{n}-nv_{n-1}=u_{n}$ for $n\in\mathbb{N}$ and $v_{0}=u_{0}$.
 In terms of formal power series we get
@@
 contradicts Proposition 2.1 because by Theorem 2.2 the formal power series $V(x)$ is rational. $\Box$
-In the original proof, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of.
+It is clear that the previous proof does not substantially use uncountable sets.
+In the original proof of the L.–W.
+theorem in [2], one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of.
 3 The proof by semiformal integration
+In this section, we present an alternative uncountable-sets-free proof of the transcendence of $\mathrm{e}$.
+It is due to this author, and it is a translation of Hilbert’s proof in the language of formal power series.
 We work in the domains of formal power series $\mathbb{R}[[x]]$ and convergent formal power series $\mathbb{R}\{x\}$; the latter are formal power series with an infinite radius of convergence.
-For $f(x)\in\mathbb{R}[[x]]$ and $n\in\mathbb{N}_{0}$ we denote by $x^{n}]f(x)$ the coefficient of $x^{n}$ in $f(x)$.
+For $f=f(x)\in\mathbb{R}[[x]]$ and $n\in\mathbb{N}_{0}$ we denote by $x^{n}]f=[x^{n}]f(x)$ the coefficient of $x^{n}$ in $f(x)$.
 For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
 $f(ax)=\sum_{n\geq 0}a^{n}a_{n}x^{n}\,.$}
-The family of operations
-$\{f(x)\mapsto f(ax)\colon\;a\in\mathbb{R}\}\,$,
-which we call constant inner products, consists of formal unary operations. For every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
+For every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
 $(af+bg)(cx)=af(cx)+bg(cx)\,\text{ and }\,(fg)(cx)=f(cx)g(cx)\,.$
-Also, $f(0x)=[x^{0}]f$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
-In the next proposition, whose straightforward proof we omit, we collect properties of the unary operations $f\mapsto f^{\prime}$ and $f\mapsto\int f$ on $\mathbb{R}[[x]]$. Recall that
+Also, $f(0x)=[x^{0}]f$. If $f(x)\in\mathbb{R}\{x\}$ then $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
+In the next proposition, whose straightforward proof we omit, we collect properties of the unary operations of the derivative $f\mapsto f^{\prime}$ and the primitive $f\mapsto\int f$ on $\mathbb{R}[[x]]$.
+Recall that
 $\int\sum_{n\geq 0}a_{n}x^{n}=\sum_{n\geq 0}\frac{a_{n}}{n+1}x^{n+1}=\sum_{n\geq 1}\frac{a_{n-1}}{n}x^{n}\,.$}
@@
   4. 4.
-    $\big(\int f\big)^{\prime}=f$ and $\int(f^{\prime})=f-[x^{0}]f$.
+    $\big(\int f\big)^{\prime}=f$ and $\int f^{\prime}=f-[x^{0}]f$.
 We define the (formal) exponential.
@@
 The formal power series
-$\mathrm{Exp}(x)=\sum_{n\geq 0}\frac{1}{n!}x^{n}=1+x+\frac{1}{2}x^{2}+\frac{1}{6}x^{3}+\dots\ \ (\in\mathbb{Q}[[x]])$}
+$\mathrm{Exp}=\mathrm{Exp}(x)=\sum_{n\geq 0}\frac{1}{n!}x^{n}=1+x+\frac{1}{2}x^{2}+\frac{1}{6}x^{3}+\dots\ \ (\in\mathbb{Q}[[x]])$}
 is called the (formal) exponential.
 Clearly, $\mathrm{Exp}\in\mathbb{R}\{x\}$, $\mathrm{Exp}^{\prime}=\mathrm{Exp}$ and $\int\mathrm{Exp}=\mathrm{Exp}-1$.
-We introduce a family of semiformal unary operations.
+We introduce a family of semiformal unary operations on $\mathbb{R}\{x\}$.
+We call them semiformal because the coefficients of the formal power series involved in these operations are obtained by limit transitions.
+In the usual formal operations with formal power series, coefficients are obtained by finite expressions.
 Definition 3.3 (shifts)
@@
 $f(x+b)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)x^{n}\;.$}
-We call these operations on $\mathbb{R}[[x]]$ semiformal because coefficients in them arise by limit transitions; in the standard formal operations they are obtained by finite expressions.
 The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent.
@@
 Clearly, $f(x+0)=f(x)$ and if $f(x)\in\mathbb{R}[x]$ then $f(x)\mapsto f(x+b)$ is the usual polynomial substitution.
-We combine the operations of constant inner product and shift by means of brackets.
+We combine the operations $f(x)\mapsto f(ax)$ and shifts by means of brackets.
 For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
-We show that shifts behave in desirable ways.
+We show that shifts behave as expected.
 Proposition 3.4
@@
 f((x+c)+d)=f(x+(c+d)).f((x+c)+d)=f(x+(c+d))\,.
-Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ and $c,d$ be as stated, and let $f(x+c)=\sum_{n\geq 0}b_{n}x^{n}$. Then
+Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ and $f(x+c)=\sum_{n\geq 0}b_{n}x^{n}$. Then
 $b_{n}=\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,f((x+c)+d)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}\big)x^{n}\,.$}
@@
 ${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}=[x^{n}]f(x+(c+d))\,.}$
-In the third line, we changed the order of summation. This is allowed because $f(x)$ is convergent. In the fourth line, we used the binomial theorem. $\Box$
+In the third line, we changed the order of summation. This is allowed because $f(x)$ is convergent. In the fifth line, we used the binomial theorem. $\Box$
 Proposition 3.5
@@
 $(f\cdot g)(x+c)=f(x+c)\cdot g(x+c)\;.$
-Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ and $c$ be as stated. Then the coefficients $x^{n}]f(x+c)$ and $x^{n}]g(x+c)$ equal to
-$\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}\,$,}
-respectively. Thus
+Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ and $g(x)=\sum_{n\geq 0}b_{n}x^{n}$. Then
+$[x^{n}]f(x+c)=\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,[x^{n}]g(x+c)=\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}\,.$}
+Thus
 ${\textstyle[x^{n}]f(x+c)g(x+c)=\sum_{k=0}^{n}\big(\sum_{m\geq k}\binom{m}{k}a_{m}c^{m-k}\sum_{l\geq n-k}\binom{l}{n-k}b_{l}c^{l+k-n}\big)}$
@@
 In the third line, we changed the order of summation. This is allowed since $f(x)$ is convergent. In the fourth line, we used Vandermonde’s identity ([7]). $\Box$
-The following semiformal exponential identity is perhaps counter-intuitive; at first we got it wrong.
+We present a curious semiformal exponential identity.
 Proposition 3.6
@@
 $\mathrm{Exp}\big((ax)+\frac{b}{a}\big)=\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$}
-Proof. Let $a$, $b$ be as stated, and let $f(x)=\mathrm{Exp}(ax)$. Then
+Proof. Let $f(x)=\mathrm{Exp}(ax)$. Then
 ${\textstyle f\big(x+\frac{b}{a}\big)}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}(a^{m}/m!)(b/a)^{m-n}\big)x^{n}}$
@@
 $\Box$
-We use this identity only for a=1a=-1.
-We introduce a family of semiformal functionals on $\mathbb{R}[[x]]$.
-Definition 3.7 (Newton integral of fps)
-Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}\in\mathbb{R}\{x\}$ and let $u,v\in\mathbb{R}$. We define the Newton integral of $f(x)$ from $u$ to $v$ as the difference
+Later we use this identity only with a=1a=-1.
+We introduce a family of semiformal functionals on $\mathbb{R}\{x\}$.
+[…]
+Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}\in\mathbb{R}\{x\}$ and $u,v\in\mathbb{R}$. We define the Newton integral of $f(x)$ from $u$ to $v$ to be the difference
 ${\textstyle\int_{u}^{v}f=\int_{u}^{v}f(x)}$ $:=$ ${\textstyle\big(\int f\big)(v)-\big(\int f\big)(u)}$
@@
 $\Box$
-The third transformation needs some work.
+To prove the third transformation, some work is needed.
 Proposition 3.10
@@
 $\int_{u}^{v}f(x+b)=\int_{u+b}^{v+b}f(x)\,.$}
-Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $u$, $v$ and $b$ be as stated. Then
+Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$. Then
 ${\textstyle\int_{u}^{v}f(x+b)=}$
@@
 This is allowed since the involved series absolutely converge.
-In the fourth line, we added
+In the fourth line, using the change of variable $l=m+1$, we added
 $0=\sum_{m\geq 0}\frac{a_{m}}{m+1}\binom{m+1}{0}b^{m+1}v^{0}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\binom{m+1}{0}b^{m+1}u^{0}\,.$}
 In the fifth line, we used the binomial theorem. $\Box$
-We introduce another family of semiformal functionals on $\mathbb{R}[[x]]$.
-Definition 3.11 (improper Newton $\int$ of fps)
+We introduce another family of semiformal functionals on $\mathbb{R}\{x\}$.
+[…]
 Let $u\in\mathbb{R}$ and $f(x)\in\mathbb{R}\{x\}$. If for every real sequence $(b_{n})$ going to $\infty$ the proper limit
@@
 $\Box$
-We need bounds on integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)\in\mathbb{Z}[x]$. We obtain them via the next bound which is based on two values of the primitive.
+We need bounds on integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)\in\mathbb{Z}[x]$. We obtain them in the next bound, which is based on two values of the primitive.
 Proposition 3.15
-For every numbers $i,k\in\mathbb{N}_{0}$,
+For every $i,k\in\mathbb{N}_{0}$,
 $\big|\int_{0}^{i}x^{k}\cdot\mathrm{Exp}(-x)\big|\leq i^{k+1}\mathrm{e}^{i}\,.$}
@@
 As we know, it is crucial that the bound is exponential and not factorial in $k$.
-We prove Euler’s identity without using uncountable sets.
+We state and prove Euler’s identity without using uncountable sets.
 Proposition 3.16 (Euler’s identity)
@@
 $\Box$
-We are ready to state Hilbert’s proof in an uncountable-sets-free way. We again assume for the contradiction that
+We are ready to phrase Hilbert’s proof in an uncountable-sets-free way. We again assume for the contradiction that
 $a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0$
@@
 for every $r\in\mathbb{N}$.
-The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as before, only now they are defined in an uncountable-sets-free way.
+The numbers $A_{r},B_{r}\in\mathbb{R}$ are as in Section 1, only now they are defined in an uncountable-sets-free way.
 We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded.
-The computation of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that now we compute semiformally.
+The computation of $B_{r}$ proceeds along the same lines as in Section 1, only now we compute semiformally in $\mathbb{R}\{x\}$.
 $B_{r}$ $=$ ${\textstyle\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{i}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)}$
@@
 has degree less than $(n+1)(r+1)$ and each coefficient in it is in absolute value at most $l^{r+1}$ where $l\in\mathbb{N}$ depends only on $n$.
-By Propositions 3.8 and 3.15 each of the above $n+1$ Newton integrals in $A_{r}$ is in absolute value at most
+By Propositions 3.8 and 3.15, each of the above $n+1$ Newton integrals in $A_{r}$ is in absolute value at most
 $(n+1)(r+1)\cdot l^{r+1}\cdot n^{(n+1)(r+1)}\cdot\mathrm{e}^{n}\,.$
 Thus $|A_{r}|\leq c^{r}$ for every $r\in\mathbb{N}$ and some constant $c\geq 1$. We get a contradiction as before. $\Box$
+This translation is more complicated than the $1\frac{1}{2}$ page proof of Hilbert at the beginning, but we eliminated uncountable sets.
+We think that the semiformal techniques used are of independent interest and will find other applications.